### v1

# TRIM: Achieving Extreme Sparsity with Targeted Row-wise Iterative Metric-driven Pruning

Abstract
Large Language Models (LLMs) present significant computational and memory challenges due to their extensive size, making pruning essential for their efficient deployment. Existing one-shot pruning methods often apply uniform sparsity constraints across layers or within each layer, resulting in suboptimal performance, especially at high sparsity ratios. This work introduces TRIM (Targeted Row-wise Iterative Metric-driven pruning), a novel approach that applies varying sparsity ratios to individual output dimensions (rows) within each layer. TRIM employs an iterative adjustment process guided by quality metrics to optimize dimension-wise sparsity allocation, focusing on reducing variance in quality retention across outputs to preserve critical information. TRIM can be seamlessly integrated with existing layer-wise pruning strategies. Our evaluations on perplexity and zero-shot tasks across diverse LLM families (Qwen2.5, LLaMA-2, and OPT) and sparsity levels demonstrate that TRIM achieves new state-of-the-art results and enhances stability. For instance, at 80% sparsity, TRIM reduces perplexity by 48% for Qwen2.5-14B and over 90% for OPT-13B compared to baseline methods. We conclude that fine-grained, dimension-wise sparsity adaptation is crucial for pushing the limits of extreme LLM compression. Code available at: https://github.com/flobk/TRIM
1 Introduction
Large language models (LLMs) have gained prominence because of their generalized problem-solving abilities (Brown et al., 2020; OpenAI et al., 2024; Zhao et al., 2025). A significant driver of this success has been the exponential scaling of model parameters (Kaplan et al., 2020; Hoffmann et al., 2022). Increasing the number of parameters has improved downstream performance and enabled emergent behaviors (Wei et al., 2022), such as in-context learning (Brown et al., 2020) and reasoning (Yao et al., 2023; Wei et al., 2023). However, this growth in model size leads to substantial memory and computational demands, which pose considerable deployment challenges (Zhou et al., 2024).
As LLM parameters scale exponentially, pruning techniques (Mozer and Smolensky, 1988; Le Cun et al., 1989; Han et al., 2015) have become increasingly important to enable efficient inference in resource-constrained settings. The goal of pruning is to reduce the number of model weights while preserving performance on downstream tasks, which helps alleviate GPU memory requirements.
Several pruning techniques have been developed to make LLMs more efficient. One major class induces sparsity during training or with fine-tuning, dynamically removing less important weights (Han et al., 2015; Evci et al., 2021; Louizos et al., 2018). While effective, these methods often require modifications to the training process and can be computationally expensive, particularly for models with billions of parameters, where maintaining optimizer states and gradients becomes a significant bottleneck (Molchanov et al., 2019; Liu et al., 2019; Hoefler et al., 2021; Gale et al., 2019).
As a result, one-shot pruning methods that reduce LLM parameters without retraining to compute weight masks have gained prominence (Han et al., 2015; Frantar and Alistarh, 2023; Sun et al., 2024). A challenge with current approaches is that they often impose predefined structural constraints, such as uniform sparsity across all network layers or outputs. This imposed structure is arbitrary and can be detrimental to pruning performance, particularly as LLMs exhibit unique weight and activation characteristics, such as prominent outlier features and highly skewed activation distributions (Kovaleva et al., 2021; Dettmers et al., 2022; Rudman et al., 2023).
Layer-wise sparsity adaptation methods aim to mitigate this by assigning distinct sparsity ratios to different layers of the model, determined through algorithmic search (Li et al., 2024; Xu et al., 2025), heuristic-based approaches (Yin et al., 2024; Sun et al., 2025), or more theoretically grounded strategies (Lu et al., 2024). While this offers more flexibility than globally uniform sparsity, these methods still apply a consistent sparsity target within each layer and can struggle under extreme sparsity levels, such as removing more than 70% of model parameters.
In this work, we propose TRIM: (Targeted Row-wise Iterative Metric-driven pruning). TRIM builds on the principles of one-shot pruning algorithms like Wanda (Sun et al., 2024), which compare and remove weights at each layer iteratively. Current methods apply a uniform sparsity constraint across all output dimensions, however, TRIM introduces a more granular approach. TRIM calculates unique sparsity ratios for individual output dimensions (corresponding to a row in the layer’s weight matrix). This fine-grained allocation is guided by the objective of reducing variance in quality retention across different outputs during the pruning process. We demonstrate that TRIM achieves state-of-the-art pruning performance at extreme sparsity levels (e.g., upwards of 80% of parameters removed).
Because TRIM adapts sparsity at the dimension-wise level, it is inherently compatible with and can augment existing layer-wise allocation strategies. TRIM can be readily integrated into importance-score-based, one-shot pruning algorithms while maintaining computational efficiency. The contributions of our work are as follows:
  1. 1.
    We introduce TRIM, the first gradient-free, one-shot pruning algorithm designed to calculate sparsity ratios on a per-dimension basis.
  2. 2.
    Using TRIM, we achieve state-of-the-art pruning results at extreme levels of sparsity reducing the perplexity of OPT-13B at 80% from 6461 with Wanda to 324 with TRIM.
  3. 3.
    We conduct a thorough empirical analysis as to why TRIM leads to superior pruning results and show that individual output dimensions differ in their 1) sensitivity to pruning and 2) criticality for down-stream performance.
Figure 1: Illustrating non-uniform, dimension-wise sparsity. On the left is Wanda, which applies the layer sparsity ratio $\mathbf{T}$ uniformly to all output dimension (rows) of the weight matrix $\mathbf{W}$. TRIM iteratively defines sparsity ratios for individual dimensions in a non-uniform way. This targeted distribution of the available sparsity budget improves local (and global) pruning quality.
2 Related Work
A large body of work has explored pruning techniques that rely on gradient information or require retraining to recover performance. Gradient-based methods such as SNIP (Lee et al., 2019), GraSP (Wang et al., 2020), and SynFlow (Tanaka et al., 2020) prune weights based on sensitivity estimates derived from gradients at initialization. Iterative methods like OBD (Le Cun et al., 1989), OBS (Hassibi et al., 1993), and movement pruning (Sanh et al., 2020) leverage gradients or Hessians during or after training to inform pruning decisions, typically requiring retraining or fine-tuning to regain lost performance (Molchanov et al., 2019; Liu et al., 2019; Frankle and Carbin, 2019).
While these methods achieve strong results, the computational burden of retraining models with billions of parameters has driven the development of efficient one-shot pruning approaches that eliminate the need for retraining. SparseGPT (Frantar and Alistarh, 2023) is the first to apply one-shot pruning to large LLMs. However, to compensate for pruning errors, it requires updating the remaining weights by calculating expensive Hessian inversions. Wanda (Sun et al., 2024) offers a more efficient alternative without updating weights. It addresses the limitations of magnitude pruning (Han et al., 2015) (which fails due to LLMs having outlier activations) by incorporating input statistics of activations. Wanda assigns a score to each weight based on both the absolute weight value and the L2 norm of a set of calibration inputs ($|W|\cdot||X||_{2}$). Further, Wanda innovates the selection of the comparison group - the subset of candidate weights from which the lowest-scoring weights are identified and removed to meet a sparsity target $T$. Wanda adopts a per-output (or row-wise) comparison, pruning the $T\%$ of lowest-scoring weights independently for each output dimension in the weight matrix. This per-output strategy departs from prior approaches, like Magnitude pruning, which utilize the entire layer as the comparison group. However, the same per-output structure that contributes to Wanda’s efficiency can also be a limitation, causing it to struggle at higher sparsity ratios compared to techniques with weight updates like SparseGPT (Frantar and Alistarh, 2023).
Methods such as Outlier Weighed Layerwise (OWL) sparsity extend Wanda by introducing layer-wise sparsity ratios (Yin et al., 2024). OWL assigns unique sparsity ratios to individual layers of a model while maintaining a global sparsity target. By assigning unique sparsity ratios to each layer, OWL avoids pruning layers that contain many outlier positions. An outlier position exists when its Wanda score, $A_{ij}$, is greater than a predefined multiple, $M$, of the mean Wanda score of that layer, $\bar{A}$. Namely, $A_{ij}$ is an outlier position if $A_{ij}>M\cdot\bar{A}$. AlphaPruning (Lu et al., 2024) offers an alternative to OWL by assigning layer-wise sparsity ratios based on the heavy-tailed shape of spectral densities in layer weight matrices. It has its roots in heavy-tailed self-regularization theory (Martin and Mahoney, 2019a, b), making it a more principled approach than OWL.
Current pruning methods, such as Wanda and its derivations OWL, and AlphaPruning, apply a uniform sparsity ratio to all output dimensions within a given layer. We demonstrate that this uniform allocation is not optimal when pruning at high levels of sparsity because different output dimensions have varying sensitivities to pruning. To address this, we propose a more granular approach by relaxing the constraint of uniform sparsity per dimension. Instead, TRIM assigns varying sparsity ratios to each output dimension $W_{i,:}$ of a weight matrix. Importantly, while the sparsity ratio can differ per output dimension, the original mechanism for selecting which weights to prune within that dimension (up to its assigned ratio) remains the same as the existing algorithm TRIM adapts.
3 TRIM - Targeted Row-wise Iterative Metric-driven pruning
Algorithm 1 Iterative Dimension-Wise Sparsity Adjustment
1:  Input: Weight matrix $W$, Input activations $X$, Target average sparsity $T$, learning rate $\alpha$
2:  Functions: Prune(), Qmetric(), QmetricDimwise()
3:  Output: Optimal dimension-wise sparsity vector $S_{best}$.
4:  Compute unpruned output: $Y\leftarrow WX$
5:  Start with uniform sparsity: $S_{i}\leftarrow T$ for $i=1,\dots,D$
6:  Track best sparsity vector: $S_{best}\leftarrow S$
7:  Track best quality score: $q_{best}\leftarrow-\infty$
8:  for $k=0$ to $K-1$ do
9:     Obtain pruned weights: $W_{pruned}\leftarrow$ Prune(W, S)
10:     Compute pruned output: $\hat{Y}\leftarrow W_{pruned}X$
11:     Calculate overall pruning quality for this iteration: $q_{k}\leftarrow\text{Qmetric}(Y,\hat{Y})$
12:     if $q_{k}>q_{best}$ then
13:        $q_{best}\leftarrow q_{k}$
14:        $S_{best}\leftarrow S$
15:     end if
16:     Calculate quality per-output: $c_{i}\leftarrow\text{QmetricDimwise}(Y_{i,:},\hat{Y}_{i,:})$ for $i=1,\dots,D$
17:     Normalize similarities to [0, 1]: $c^{\prime}_{i}\leftarrow\frac{c_{i}-\min_{j}c_{j}}{\max_{j}c_{j}-\min_{j}c_{j}+\epsilon}$ for $i=1,\dots,D$
18:     Apply learning rate: $\delta_{i}\leftarrow\alpha c^{\prime}_{i}$ for $i=1,\dots,D$
19:     Recenter mean: $S_{i}\leftarrow\delta_{i}-\frac{1}{D}\sum_{j=1}^{D}\delta_{j}+T$ for $i=1,\dots,D$
20:  end for
21:  Return optimal sparsity allocation: Return $S_{best}$
Figure 1 illustrates the core concept of TRIM. Let $W\in\mathbb{R}^{D\times N}$ be the weight matrix of a layer, where $D$ is the number of output dimensions (rows) and $N$ is the number of input dimensions (columns). Each row $W_{i,:}$ corresponds to the weights associated with the $i$-th output dimension. We define a dimension-wise sparsity vector $S=[S_{1},S_{2},\dots,S_{D}]$, where each element $S_{i}\in[0,1]$ specifies the target sparsity ratio for the $i$-th output dimension $W_{i,:}$. This means that for each row $W_{i,:}$, $S_{i}\cdot N$ of its weights will be pruned.
While the individual sparsity ratios $S_{i}$ can vary, they must collectively satisfy the target sparsity $T$ for the layer. This is enforced by ensuring their average equals $T$:
$\frac{1}{D}\sum_{i=1}^{D}S_{i}=T$ (1)
Setting $S_{i}=T$ for all $i\in\{1,\dots,D\}$ reverts to the standard pruning approach where sparsity is uniform across output dimensions.
The core challenge is to determine an effective allocation strategy for the dimension-wise sparsity vector, $S$, that optimizes performance while adhering to Equation 1. Below, we detail our method for finding $S$, given a layer’s weight matrix, $W$, and $L$ sample input activations $X\in\mathbb{R}^{N\times L}$.
Calculating Dimension-Wise Sparsity Vectors.
To find a suitable dimension-wise sparsity vector $S$, we employ an iterative adjustment process, detailed in Algorithm 1. The process begins by computing the unpruned output, $Y\leftarrow WX$, (Line 4) with $Y\in\mathbb{R}^{D\times L}$ and initializing $S$ uniformly, such that $S_{i}=T$ for all output dimensions $i=1,\dots,D$ (Line 5). This initial $S$ also serves as the first candidate for the best sparsity vector, $S_{best}$.
The algorithm then iterates $K$ times (Lines 8-20). In each iteration $k$:
  1. 1.
    The weight matrix $W$ is pruned using the current sparsity vector to obtain $W_{pruned}$ (Line 9). This means for each row $W_{i,:}$, $S_{i}\cdot N$ weights are removed based on the chosen pruning criterion (e.g., lowest Wanda scores).
  2. 2.
    The pruned output $\hat{Y}\leftarrow W_{pruned}X$ is computed (Line 10).
  3. 3.
    The overall quality $q_{k}\leftarrow\text{Qmetric}(Y,\hat{Y})$ of the iteration is calculated (Line 11). If $q_{k}$ is better than the best quality found so far ($q_{best}$), then $q_{best}$ is updated, and the current $S$ is stored as $S_{best}$ (Lines 12-14).
  4. 4.
    The quality $c_{i}\leftarrow\text{QmetricDimwise}(Y_{i,:},\hat{Y}_{i,:})$ for each individual output dimension is calculated (Line 15).
  5. 5.
    The sparsity targets $S_{i}$ are adjusted based on these per-dimension quality scores. First, the scores $c_{i}$ are normalized to a [0, 1] range (Line 16). An update term $\delta_{i}\leftarrow\alpha c^{\prime}_{i}$ is calculated using a learning rate $\alpha$ (Line 17). This mechanism aims to increase the sparsity target $S_{i}$ for dimensions exhibiting higher quality (less degradation, higher $c^{\prime}_{i}$) and decrease it for dimensions with lower quality, thereby pruning less sensitive dimensions more aggressively.
  6. 6.
    Finally, the adjusted sparsity vector is re-centered to ensure its average value remains equal to the overall target layer sparsity $T$ (Line 18) in order to satisfy Equation 1.
After $K$ iterations, the algorithm returns $S_{best}$ (Line 21), which represents the dimension-wise sparsity allocation that yields the highest overall quality $q_{best}$ during the iterative search. $S_{best}$ is then used to perform the final pruning of the layer.
To find the most effective sparsity distribution, we first explore positive learning rates, $\alpha$. We start with a small $\alpha$, run the iterative adjustment process (Algorithm 1), and record $q_{best}$. We then incrementally increase $\alpha$ and repeat the process, stopping when a larger $\alpha$ no longer improves $q_{best}$. However, if a positive $\alpha$ does not yield a quality $q_{best}$ superior to the initial baseline (uniform sparsity), we use a negative learning rate. This allows for better generalization across different model activation spaces.
We use cosine similarity as our quality metric and refer the reader to Appendix A for the evaluation of other metrics. Additionally, Appendix B presents our analysis of the computational overhead of this iterative procedure. Our findings show that calculating the sparsity vectors for all layers of the Qwen2.5-14B model adds only a small delay, taking between 27.76 and 59.88 seconds based on the number of calibration samples used.
Table 1: Perplexity results on WikiText validation set for various models pruned at 70% and 80% sparsity. The $\Delta$ column shows the average perplexity reduction across the tested models achieved by adding TRIM. Lower is better.
Method OPT LLaMA-2 Qwen2.5 6.7B 13B 7B 13B 3B 7B 14B 32B 72B $\Delta$ 0% - 10.86 10.13 5.47 4.88 8.03 6.85 5.29 5.018 3.851 - 70% Alpha 63.53 820.42 31.32 15.19 164.39 68.93 61.59 15.70 11.66 -25.3% +TRIM 30.80 35.16 29.00 14.70 128.44 62.22 46.43 13.94 11.41 OWL 40.30 28.92 21.20 16.02 83.98 44.60 36.46 13.06 11.25 -4.6% +TRIM 38.21 28.21 20.10 14.39 83.39 43.97 33.21 12.28 11.06 80% Alpha 5357.60 4726.09 1558.41 143.19 9872.21 728.82 1040.91 469.52 70.26 -41.6% +TRIM 656.26 557.30 1527.01 108.70 1685.51 1037.11 435.84 133.94 68.64 OWL 8509.51 6461.43 406.20 225.04 768.69 245.32 348.48 73.44 61.59 -34.9% +TRIM 492.47 324.14 437.88 154.83 504.89 233.80 180.67 66.95 58.15
Figure 2: Perplexity progression from 70% to 80% sparsity. TRIM-enhanced OWL shows delayed and slower perplexity growth, effectively extending the usable sparsity frontier.
4 Experiments
Models and Evaluation.
We evaluate TRIM on Qwen2.5 3B/7B/14B/32B/72B (Qwen et al., 2025), LLaMA-2 7B/13B (Touvron et al., 2023), and OPT 7B/13B (Zhang et al., 2022). Our evaluation adheres to established methodologies for pruning LLMs (Sun et al., 2024; Frantar and Alistarh, 2023). We evaluate language modeling proficiency by calculating perplexity on the reserved WikiText validation set (Merity et al., 2016) and assess zero-shot performance across a range of down-stream tasks. We evaluate on BoolQ (Clark et al., 2019), RTE (Wang et al., 2019), HellaSwag (Zellers et al., 2019), WinoGrande (Sakaguchi et al., 2019), ARC Easy and Challenge (Clark et al., 2018), and OpenbookQA (Mihaylov et al., 2018), using the framework by Gao et al. (Mihaylov et al., 2018).
Baselines.
We evaluate Wanda (Sun et al., 2024) based OWL (Yin et al., 2024) and AlphaPruning (Lu et al., 2024) with and without TRIM. For all runs, we use the same number of calibration samples, each with a sequence length of 2048 tokens, randomly selected from the first shard of the C4 dataset (Raffel et al., 2023). More details and hyperparameters for the layerwise methods can be found in 11.
4.1 Main results
We evaluate pruning at high sparsity ratios of 70% and 80%. For all methods, we focus on pruning the linear layers, excluding the initial embedding layer and the final classification head.
Language Modeling.
Table 1 shows perplexity results for all models used in this study. Integrating TRIM with Wanda, OWL and AlphaPruning consistently results in lower perplexity values than these baselines without TRIM. At 70% and 80% sparsity respectively, TRIM reduces perplexity for LLaMA-2-13B by 10% and 31%, and for Qwen2.5-14B by 9% and 48%. The effectiveness of TRIM is especially pronounced for OPT-13B at 80% sparsity, where TRIM decreases OWL pruning perplexity from 6461.43 to 324.14.
TRIM also improves the robustness and reliability of pruning outcomes. Baseline methods can occasionally lead to counterintuitive scaling results. For example, pruning Qwen2.5-14B at 80% sparsity has a higher perplexity than its 7B counterpart. Baseline methods fail to effectively prune OPT models at 80% sparsity, resulting in very high perplexity. TRIM delivers substantial improvements for OPT stabilizing performance at high sparsity levels. These consistent and robust improvements across all model families, sizes, and sparsity levels underscore the critical role of effective dimension-wise sparsity allocation and highlight TRIM’s efficacy in achieving it.
Zero-Shot Downstream Tasks.
The benefits of TRIM extend beyond perplexity improvements to generalization on downstream tasks. As evidenced by the average zero-shot accuracies presented in Table 2, integrating TRIM with layer-wise pruning strategies yields superior performance compared to baseline methods without TRIM. This performance advantage holds across a wide range of models and sparsity levels. The complete result tables can be found in Appendices 12 and 13.
Table 2: Average zero-shot accuracies (%) for models pruned at 70% and 80% sparsity. $\Delta$ shows the average percentage increase achieved by adding TRIM. Higher is better.
Method OPT LLaMA-2 Qwen2.5 6.7B 13B 7B 13B 3B 7B 14B 32B 72B $\Delta$ 0% - 51.52 52.60 59.73 63.02 61.13 65.82 68.12 68.09 70.13 - 70% Alpha 38.42 39.35 42.01 47.39 35.43 39.65 39.92 52.98 60.80 +1.14% +TRIM 42.27 42.27 41.60 47.92 36.75 39.75 40.73 53.15 61.76 OWL 40.67 43.60 43.30 49.01 36.53 39.22 40.92 54.56 61.30 +0.42% +TRIM 41.16 43.12 43.36 49.54 36.65 39.82 41.68 55.04 62.47 80% Alpha 32.70 35.13 31.94 36.71 32.77 32.74 35.32 36.06 40.32 +0.65% +TRIM 35.67 36.56 31.99 36.73 33.16 32.98 34.40 37.44 40.59 OWL 32.61 35.22 32.60 36.23 35.42 32.31 33.70 37.35 41.19 +0.46% +TRIM 36.30 35.47 32.79 36.87 34.82 32.64 33.32 37.48 41.02
4.2 Generalization To Other Pruning Metrics.
Although we primarily evaluate TRIM with the Wanda (Sun et al., 2024) metric, the dimension-wise sparsity and per-output optimization of TRIM can be applied to other pruning metrics. To test the generalization capabilities of TRIM, we evaluate using the SparseGPT pruning metric (Frantar and Alistarh, 2023), which assigns a score to each weight based on the squared value of the weight divided by the corresponding diagonal element of the inverse Hessian approximation $((\mathbf{X}^{T}\mathbf{X})^{-1})_{ii}$, where $\mathbf{X}$ are the input activations to the layer. SparseGPT groups weights in blocks of 128 inputs (columns of the weight matrix), as opposed to the per-output (a single row) comparison group of Wanda. We omit the weight updates in SparseGPT to focus on the pruning configuration and the potential improvement capabilities of TRIM. We also evaluate the magnitude pruning metric (Han et al., 2015), which uses the absolute value of the weights as a score, $|\mathbf{W_{ij}}|$, and selects pruning candidates by comparing weights within the entire layer. Table 3 shows that the per-output pruning configuration with TRIM performs best, confirming that TRIM generalizes across different metrics.
Table 3: Wikitext validation perplexity of LLaMA-2-13B and Qwen2.5-14B pruned at 70% sparsity with OWL layer ratios. TRIM improves upon the per-output configuration and generalizes to other pruning metrics besides Wanda. Underscored results indicate the default pruning configuration.
Model Method Pruning Metric (layer) (input, 128) (1, output) (1, output) +TRIM LLaMA-2-13B Magnitude $|\mathbf{W_{ij}}|$ 59.86 49.25 23.48 19.94 Wanda $|\mathbf{W_{ij}}|\cdot||\mathbf{X_{j}}||$ 20.36 24.64 16.02 14.39 SparseGPT $[|\mathbf{W}|^{2}/\text{diag}(\mathbf{H^{-1}})]_{ij}$ 19.75 19.23 15.19 14.11 Qwen2.5-14B Magnitude $|\mathbf{W_{ij}}|$ 2e3 2e3 3e4 2e3 Wanda $|\mathbf{W_{ij}}|\cdot||\mathbf{X_{j}}||$ 40.92 43.54 36.46 33.21 SparseGPT $[|\mathbf{W}|^{2}/\text{diag}(\mathbf{H^{-1}})]_{ij}$ 24.71 27.93 24.18 23.50
5 Analysis
In this section, we highlight three key empirical observations that explain TRIM’s effectiveness: 1) output dimensions exhibit varying sensitivity to pruning; 2) post-pruning quality degrades as sparsity levels increase; and 3) some output dimensions are critical to model performance, while others contribute little to model performance.
1. Output dimensions exhibit different sensitivity to pruning.
An output dimension’s pruning sensitivity is closely tied to its importance score concentration. We examine the concentration of Wanda scores ($|W_{i,:}|\cdot||X||_{2}$) for each output dimension using Gini coefficients. Figure 3 displays histograms of these per-dimension Gini coefficients for the first, middle and last layer of Qwen2.5-14B. Figure 3 demonstrates that some dimensions rely on a few key weights while others distribute importance more broadly. These trends are consistent across all models and model layers.
(a) Layer 1
(b) Layer 24
(c) Layer 48
Figure 3: Histograms of Gini coefficients for the Wanda pruning metric across output dimensions, in different layers of Qwen2.5-14B (K-proj). A high Gini coefficient indicates that the signal for that dimension is concentrated in a few weights.
The heterogeneity of importance concentration translates directly into differing pruning sensitivities across output dimensions. To demonstrate this, we increase the uniform sparsity ratio applied to a layer and measure the resulting quality degradation for each individual output dimension. We quantify the quality using the cosine similarity $c_{i}=\text{CosineSimilarity}(Y_{i,:},\hat{Y}_{i,:})$ between the original output $Y_{i,:}$ and the pruned output $\hat{Y}_{i,:}$ for that dimension, given a set of sample inputs $X$. Figure 4 plots these cosine similarities $c_{i}$ against increasing levels of sparsity $T$. The rate at which quality degrades differs significantly across dimensions. Some dimensions maintain high similarity even at high sparsity levels (indicating resilience to pruning), while others experience a rapid drop even at lower sparsity (indicating sensitivity to pruning). This result confirms that output dimensions are indeed differently “difficult” to prune and that this difficulty is linked to their importance concentration.
Figure 4: Cosine similarity for all output dimensions (Qwen2.5-14B, layer-24, k-proj). Increasing sparsity shows varying pruning sensitivity.
2. Quality degradation accelerates with increasing sparsity.
Pruning can remove redundant weights with minimal impact at low sparsity levels due to the overparameterization of LLMs. However, as sparsity rises, pruning inevitably affects weights carrying more essential activation signals, leading to a faster decay in performance. This non-linear decline is evident in the average trend of quality degradation on the dimension-level (Figure 4) or in downstream metrics like the exponential increase in perplexity at high sparsity rates (Figure 2).
3. Output dimensions differ in their impact on overall performance.
Beyond the differences in their sensitivity to pruning, output dimensions vary significantly in their importance for the model’s performance. To illustrate this, we conduct an experiment with Qwen2.5-3B, where we fully eliminate exactly one output dimension per layer, resulting in 0.03% total sparsity (Table 5). Removing the dimension with the smallest L2-norm only increases perplexity by 0.16 from the unpruned baseline, similar to choosing a random dimension (+0.15). However, pruning the dimension with the highest norm increases perplexity to 273.10. The importance of a dimension can also be linked to the concentration of outlier positions present in it. We define an outlier position using the OWL criterion, where the Wanda metric score $A_{ij}$ is at least a multiple of the layer mean $\bar{A}$. We take the top 10% of dimensions with the largest number of outliers (i.e., “outlier-dense” dimensions) and prune them with 90% sparsity, resulting in 9% total model sparsity. Note that pruning this way retains all of the outlier positions since they are always the last weights to get pruned. Table 5 shows that pruning outlier-dense dimensions significantly impacts performance compared to randomly pruning an equivalent number of dimensions with the same sparsity. This result highlights two key points.
Table 4: Wikitext validation perplexity for Qwen2.5-3B. Exactly one dimension gets removed.
Method Perplexity
Baseline 8.03
$D$ w. min norm 8.19
$D$ w. max norm 273.10
Random $D$ 8.18
Table 5: Wikitext validation perplexity. 10% of dimensions $D$ are heavily pruned (90%). Outlier-dense dimensions are important, however differently so across model families.
Model Baseline Random Outlier-dense $D$
Qwen2.5-7B 6.85 9.17 26.39
Qwen2.5-14B 5.29 7.70 26.16
LLaMA2-7B 5.47 7.27 255.06
LLaMA2-13B 4.88 6.23 74.27
First, outlier-dense dimensions are important for model performance. Second, although outlier positions in an outlier-dense dimension may appear to capture most of the signal, our findings show that non-outlier weights also carry significant informational value, limiting the effectiveness of aggressive pruning. Further, the extent of this impact differs between model families. For example, LLaMA models exhibit a more pronounced degradation than Qwen, indicating that outlier-dense dimensions are more important for some models than for others. Together, these findings highlight that output dimensions are not uniformly important and that simple heuristics, such as L2-Norm or outlier characteristics, can identify dimensions critical for preserving model capabilities.
6 Discussion
Targeted sparsity allocation has proven successful for enhancing the performance of pruned LLMs, but has so far only been attempted at the layer level (Yin et al., 2024; Lu et al., 2024; Sun et al., 2025; Xu et al., 2025). This work extends this paradigm to individual output dimensions within each layer. This is particularly promising, and indeed necessary, when aiming for high sparsity levels. As our findings indicate (Observation 5), post-pruning quality tends to degrade at an accelerated rate with increasing sparsity, making indiscriminate or overly coarse pruning strategies increasingly detrimental. The core premise of TRIM and the concept of dimension-wise sparsity allocation is that there exists significant heterogeneity to exploit at the dimension level (Observation 5 and 5). However, moving to such a fine-grained level of control introduces challenges: the search space for optimal sparsity configurations becomes substantially larger, and the feedback signals (e.g., quality metrics per dimension) might become less reliable or more noisy compared to layer-level aggregates.
Figure 5: LLaMA-2-13B (left) with high Gini coefficients (concentrated outliers) uses more negative learning rates than Qwen2.5-14B (right) with low Ginis (uniform outliers). Negative LRs preserve critical, outlier-dense dimensions with high initial quality. Both plots show the gate-projection.
TRIM is designed to navigate this complexity through two primary ideas. First, it recognizes that, in many cases, reducing the variance in post-pruning quality degradation across dimensions is beneficial for overall layer stability. This is pursued via a positive learning rate in its iterative refinement process (Algorithm 1), aiming to prevent any single dimension from experiencing significant quality loss. Second, TRIM acknowledges that a simple quality equalization strategy is not universally optimal. Some output dimensions possess both high initial post-pruning quality and are critically important for the model’s function (e.g., outlier-dense dimensions as shown in Table 5). Forcing these dimensions to match the potentially lower quality of less important and more sensitive ones through aggressive variance reduction would be counterproductive. This leads to TRIM’s adaptive mechanism, which involves the monitoring of layer quality and the learning rate adaptation. When beneficial, TRIM can even employ a negative learning rate, effectively increasing quality variance. LLaMA models, for instance, often exhibit strong outlier-dense dimensions (reflected in high Gini coefficients for outlier concentration, Figure 5) and, as a result, TRIM frequently selects a negative learning rate for these models (approximately 40% of the time, compared to 10% for other models). While TRIM’s flexible learning rate and best-state selection address this inherent tension in optimization direction to a significant extent, it remains a limitation and exploring alternative adjustment mechanisms or more direct methods to balance these competing objectives could be promising for future research.
7 Conclusion
In this paper, we introduce TRIM (Targeted Row-wise Iterative Metric-driven pruning), a novel approach that, for the first time, extends targeted sparsity allocation not across layers but within each layer by introducing the concept of dimension-wise sparsity. TRIM enables stable performance at high target sparsity by iteratively refining a sparsity ratio for each output dimension of a layer. Using TRIM, we reduce the perplexity by 48% for Qwen2.5-14B and over 90% for OPT-13B compared to existing state-of-the-art methods when pruning at 80% sparsity. TRIM consistently improves pruning results when paired with layer-wise sparsity allocation techniques across OPT, LLaMA, and Qwen2.5 without extensive computational overhead. By enabling more granular control, TRIM offers a more effective and reliable path towards pruning LLMs at high sparsity targets, contributing to a deeper understanding of fine-grained sparsity allocation and paving the way for future advancements in model compression.
Appendix A Quality Metric
TRIM utilizes two distinct quality evaluations to guide the pruning process. The first, layer-wise quality evaluation, assesses the pruning quality for an entire layer. We conduct an ablation study across different sparsity levels, model sizes, and model families, with results presented in Table 6. The metrics considered are Mean Squared Error (MSE), Cosine Flat (which flattens the two-dimensional output vectors before calculating cosine similarity), and Cosine Sample (which calculates cosine similarity along the sample axis and then averages the results). Our findings indicate that both Cosine metrics offer a robust balance between accuracy and reliability across diverse pruning scenarios.
Table 6: Perplexity results for different layer quality metrics.
Model Sparsity Cosim Flat Cosim Sample MSE
Qwen2.5-3B 60% 19.35 19.38 19.42
OPT-6.7B 70% 32.51 32.75 29.62
OPT-6.7B 80% 602.57 590.12 1970.02
Qwen2.5-14B 80% 184.36 170.23 189.66
LLaMA-2-13B 80% 162.33 173.09 142.81
The second evaluation, dimension-wise quality assessment, measures the post-pruning quality for each individual output dimension. For this assessment, we investigate Cosine Similarity, Mean Squared Error (MSE), and Peak Signal-to-Noise Ratio (PSNR). As shown in Table 7, while MSE can outperform Cosine Similarity in specific cases (e.g., for LLaMA models), Cosine Similarity generally demonstrates greater reliability.
Table 7: Perplexity results for different dimension-wise quality metrics.
Model Sparsity Cosine Similarity PSNR MSE
Qwen2.5-3B 60% 19.35 19.33 19.26
OPT-6.7B 70% 32.51 38.64 44.72
OPT-6.7B 80% 602.57 6834.343 4405.93
Qwen2.5-14B 80% 184.36 216.87 276.01
LLaMA-2-13B 80% 162.34 226.88 137.57
These results demonstrate that: a) Cosine Similarity, as the default configuration for both layer-wise and dimension-wise quality assessments, proves to be a reliable choice, b) TRIM’s performance can be further optimized in specific pruning scenarios by performing a hyperparameter sweep to identify the most suitable quality metric, and c) the choice of dimension-wise quality metric is typically more important than the layer-wise metric for overall pruning performance.
Appendix B Runtime Overhead
One-shot pruning methods like Wanda (Sun et al., 2024) are designed for efficient pruning of models with billions of parameters. Any extension to these methods should maintain this computational efficiency. TRIM, while iterative in nature, is designed around that and performs most computations in parallel on the GPU.
To empirically validate this claim, we prune the Qwen2.5-14B model using one NVIDIA A100-40GB and report the time in Table 8. TRIM adds little overhead compared to the duration of the entire Wanda based pruning function (8% runtime increase for 32 samples, and 1.8% increase for 512 samples).
Table 8: Runtime overhead of TRIM with different numbers of calibration samples.
N samples 32 64 128 256 512
Wikitext Perplexity 34.70 33.06 33.49 32.72 33.21
TRIM overhead (s) 27.76 31.78 35.96 42.45 59.88
Total Time (s) 375 562 888 1904 3325
Appendix C Robustness Analysis
To evaluate the robustness, we run TRIM five times with different random seeds and report the results in Table 9. TRIM slightly increases the standard deviation compared to the OWL baseline method.
Table 9: Robustness analysis of TRIM with different random seeds on Qwen2.5-14B and LLaMA-2-7B at 70% sparsity. Wikitext validation perplexity, lower is better.
Model Method Seed 1 Seed 2 Seed 3 Seed 4 Seed 5 Mean Std
Qwen2.5-14B OWL 37.63 37.43 36.33 37.71 36.87 37.20 0.52
Qwen2.5-14B +TRIM 33.64 33.50 33.61 33.16 34.95 33.78 0.61
LLaMA-2-7B OWL 21.53 21.39 21.32 21.58 21.60 21.48 0.11
LLaMA-2-7B +TRIM 20.44 20.37 19.95 20.40 20.73 20.38 0.25
Appendix D Implementation Details
Sparsity Cutoff.
The limited amount of calibration data can push the sparsity for a dimension into a local minimum, sometimes completely eliminating it. To prevent this overfitting from happening, we limit the sparsity that any dimension can experience to at most 95%. This helps TRIM achieve more stable results.
Input Recalculation.
Additionally, we give the option for input vector recalculation. One-shot pruning methods like Wanda prune block by block, and in each block layer by layer. However, when the Q,K,V-projections are pruned the activation statistic for subsequent layers (Out-projection, or the MLP block) also changes. Therefore, we include an option to recalculate the input vectors used by TRIM. Note that this does not include the Wanda metric itself, which ensures fair comparisons.
This recalculation makes pruning quality slightly better (see Table 10) and results in increased inference time. We recommend enabling this option for achieving the best results, and disabling it during hyperparameter tuning or developing.
Table 10: Effect of input recalculation on perplexity for Qwen2.5-14B at different sparsity levels.
Sparsity Baseline Without Recalc With Recalc
70% 36.46 33.85 33.21
80% 348.48 186.32 180.67
Appendix E Hyperparameters
We publish the hyperparameters used for the layer-wise sparsity allocation methods. For OWL (Yin et al., 2024) we searched in the intervals $\lambda\in\{0.02,0.05,0.08,0.12,0.15,0.2\}$ and $\mathbf{M}\in\{3,5,7,10\}$. For AlphaPruning (Lu et al., 2024) we searched in the interval $\tau\in\{0.1,0.2,0.3,0.4\}$. All sweeps were done at 70% sparsity with one seed and Wanda as the underlying (Sun et al., 2024) pruning metric.
Table 11: Hyperparameters for the layerwise sparsity allocation methods that produce the results in this paper.
Model OWL (M, $\lambda$) AlphaPruning ($\tau$)
OPT-6.7B (10, 12%) 0.1
OPT-13B (10, 5%) 0.1
LLaMA-2-7B (5, 20%) 0.3
LLaMA-2-13B (5, 15%) 0.3
Qwen2.5-3B (5, 12%) 0.1
Qwen2.5-7B (5, 12%) 0.1
Qwen2.5-14B (3, 20%) 0.3
Qwen2.5-32B (5, 15%) 0.3
Qwen2.5-72B (7, 8%) 0.1
Appendix F Task Results
Table 12: Accuracies (%) for 7 zero-shot tasks for various models with unstructured 70% sparsity and based on Wanda. Higher is better. We add TRIM to non-uniform layer-wise allocation methods OWL and AlphaPruning.
Model Method BoolQ RTE Hella- Wino- ARC-e ARC-c OBQA Mean Swag Grande OPT-6.7B Dense 66.09 55.23 50.48 65.19 65.61 30.46 27.60 51.52 Alpha 62.14 55.96 28.99 52.49 40.07 17.06 12.20 38.41 +TRIM 62.17 53.43 35.72 54.54 48.86 22.35 18.80 42.27 OWL 62.17 51.99 32.90 53.35 47.10 19.96 17.20 40.67 +TRIM 62.20 52.71 33.13 52.80 48.27 20.82 18.20 41.16 OPT-13B Dense 65.87 57.76 52.46 65.19 67.09 32.85 27.00 52.60 Alpha 54.68 58.12 31.20 54.14 37.50 22.18 17.60 39.35 +TRIM 62.29 53.07 37.37 57.30 46.25 23.21 16.40 42.27 OWL 62.23 57.40 37.65 58.01 47.35 25.34 17.20 43.60 +TRIM 62.26 54.51 37.33 58.56 48.23 23.72 17.20 43.12 LLaMA-2-7B Dense 77.71 62.82 57.19 69.22 76.35 43.43 31.40 59.73 Alpha 62.26 52.71 34.74 61.64 43.22 22.10 17.40 42.01 +TRIM 62.45 53.07 34.74 58.56 42.97 22.18 17.20 41.60 OWL 62.35 52.71 36.91 60.93 48.11 22.87 19.20 43.30 +TRIM 62.39 52.71 37.15 60.22 47.81 23.46 19.80 43.36 LLaMA-2-13B Dense 80.55 65.34 60.05 72.14 79.42 48.46 35.20 63.02 Alpha 62.91 52.71 40.53 66.69 56.27 29.61 23.00 47.39 +TRIM 65.41 53.07 40.79 66.46 57.28 29.44 23.00 47.92 OWL 70.43 52.71 41.27 65.90 58.96 29.61 24.20 49.01 +TRIM 71.62 52.71 41.82 67.48 58.75 29.61 24.80 49.54 Qwen2.5-3B Dense 77.34 75.45 54.95 68.35 77.36 44.88 29.60 61.13 Alpha 53.91 53.43 27.72 50.04 33.33 18.17 11.40 35.43 +TRIM 61.71 52.71 27.68 50.20 34.47 17.66 12.80 36.75 OWL 62.17 52.71 27.97 49.96 33.42 17.49 12.00 36.53 +TRIM 62.17 52.71 28.00 48.70 36.07 16.89 12.00 36.65 Qwen2.5-7B Dense 84.68 81.59 60.01 73.01 80.47 47.78 33.20 65.82 Alpha 62.29 52.71 30.42 54.78 44.23 19.11 14.00 39.65 +TRIM 62.08 52.71 30.76 52.25 45.83 19.62 15.00 39.75 OWL 62.14 52.71 30.35 52.33 45.08 17.92 14.00 39.22 +TRIM 62.20 52.71 30.63 53.75 45.92 19.54 14.00 39.82 Qwen2.5-14B Dense 85.23 80.14 63.36 75.30 82.45 55.97 34.40 68.12 Alpha 62.17 52.71 29.86 54.46 47.09 18.94 14.20 39.92 +TRIM 62.20 52.71 30.66 55.56 48.57 20.22 15.20 40.73 OWL 62.17 52.07 31.82 57.14 48.36 19.28 15.60 40.92 +TRIM 62.17 52.71 32.54 58.88 50.42 19.45 15.60 41.68 Qwen2.5-32B Dense 87.12 81.59 64.98 75.22 80.85 53.07 33.80 68.09 Alpha 68.93 64.98 43.43 69.69 66.58 33.02 24.20 52.98 +TRIM 70.24 62.09 44.19 69.38 67.55 33.79 24.80 53.15 OWL 71.38 62.81 44.16 70.09 71.21 36.43 25.80 54.55 +TRIM 72.57 62.45 45.15 70.64 72.35 36.95 25.20 55.04 Qwen2.5-72B Dense 89.24 77.26 67.58 77.58 85.06 58.36 35.80 70.12 Alpha 77.43 79.06 49.83 74.35 74.12 43.00 27.80 60.80 +TRIM 80.31 80.87 50.09 74.59 74.66 43.17 28.60 61.76 OWL 76.61 78.34 50.77 74.43 76.26 44.88 27.80 61.30 +TRIM 81.80 80.50 50.92 75.06 76.26 44.11 28.60 62.47
Table 13: Accuracies (%) for 7 zero-shot tasks for various models with unstructured 80% sparsity and based on Wanda. Higher is better. We add TRIM to non-uniform layer-wise allocation methods OWL and AlphaPruning.
Model Method BoolQ RTE Hella- Wino- ARC-e ARC-c OBQA Mean Swag Grande OPT-6.7B Dense 66.09 55.23 50.48 65.19 65.61 30.46 27.60 51.52 Alpha 41.04 52.71 25.93 50.20 27.44 19.97 11.60 32.70 +TRIM 60.70 50.18 26.57 51.38 31.19 18.69 11.00 35.67 OWL 38.62 51.99 25.90 50.75 25.97 20.65 14.40 32.61 +TRIM 61.93 52.71 26.86 51.38 33.00 17.41 10.80 36.30 OPT-13B Dense 65.87 57.76 52.46 65.19 67.09 32.85 27.00 52.60 Alpha 60.83 51.26 25.83 49.49 26.94 19.96 11.60 35.13 +TRIM 61.87 52.71 26.75 52.25 29.80 19.54 13.00 36.56 OWL 56.39 52.71 26.10 51.14 26.89 20.90 12.40 35.22 +TRIM 58.01 52.71 26.99 50.67 30.35 17.75 11.80 35.47 LLaMA-2-7B Dense 77.71 62.82 57.19 69.22 76.35 43.43 31.40 59.73 Alpha 37.89 52.71 26.20 47.28 26.73 20.14 12.60 31.94 +TRIM 37.89 52.71 26.18 47.04 26.56 20.56 13.00 31.99 OWL 38.19 52.71 26.64 50.36 27.82 18.69 13.80 32.60 +TRIM 38.26 52.71 26.50 51.38 27.86 19.45 13.40 32.79 LLaMA-2-13B Dense 80.55 65.34 60.05 72.14 79.42 48.46 35.20 63.02 Alpha 62.05 52.71 27.74 52.72 30.68 19.28 11.80 36.71 +TRIM 62.17 52.71 27.90 51.62 30.85 19.88 12.00 36.73 OWL 61.41 52.71 27.49 50.75 30.35 19.11 11.80 36.23 +TRIM 62.08 52.71 28.06 53.04 30.64 19.37 12.20 36.87 Qwen2.5-3B Dense 77.34 75.45 54.95 68.35 77.36 44.88 29.60 61.13 Alpha 37.77 55.23 26.12 50.04 27.06 19.80 13.40 32.77 +TRIM 44.28 49.46 26.47 48.54 28.58 20.56 14.20 33.16 OWL 56.73 52.71 26.78 51.22 30.30 19.03 11.20 35.42 +TRIM 54.10 52.71 26.81 49.25 29.71 18.17 13.00 34.82 Qwen2.5-7B Dense 84.68 81.59 60.01 73.01 80.47 47.78 33.20 65.82 Alpha 37.92 52.71 26.91 49.88 28.37 19.03 14.40 32.74 +TRIM 37.77 53.06 27.21 49.96 30.01 18.86 14.00 32.98 OWL 37.89 52.71 27.14 47.28 30.51 17.83 12.80 32.31 +TRIM 38.29 52.71 27.30 47.83 30.09 18.09 14.20 32.64 Qwen2.5-14B Dense 85.23 80.14 63.36 75.30 82.45 55.97 34.40 68.12 Alpha 59.17 52.71 26.94 48.93 30.43 16.64 12.40 35.32 +TRIM 50.67 52.71 26.86 50.12 30.89 16.55 13.00 34.40 OWL 44.59 52.71 27.27 49.17 31.94 16.64 13.60 33.70 +TRIM 42.08 52.71 27.08 49.25 32.15 16.98 13.00 33.32 Qwen2.5-32B Dense 87.12 81.59 64.98 75.22 80.85 53.07 33.80 68.09 Alpha 59.14 52.71 27.61 48.22 35.19 17.15 12.40 36.06 +TRIM 61.83 53.43 28.79 49.49 37.46 17.49 13.60 37.44 OWL 62.08 52.71 28.49 49.57 38.38 17.24 13.00 37.35 +TRIM 62.20 52.71 28.35 50.20 39.18 16.30 13.40 37.48 Qwen2.5-72B Dense 89.24 77.26 67.58 77.58 85.06 58.36 35.80 70.12 Alpha 62.17 52.71 32.14 58.41 41.12 21.50 14.20 40.32 +TRIM 62.17 52.71 32.50 58.72 41.24 21.16 15.60 40.59 OWL 62.17 52.71 32.63 58.88 43.56 21.67 16.20 41.12 +TRIM 62.17 52.71 32.69 57.85 44.28 21.42 16.00 41.02
NeurIPS Paper Checklist
  1. 1.
    Claims
  2. Question: Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope?
  3. Answer: [Yes]
  4. Justification: The abstract and introduction clearly state the contributions. The contributions are made based on the presentation of our method (3), the results in 4, and the analysis in 5.
  5. Guidelines:
    • The answer NA means that the abstract and introduction do not include the claims made in the paper.
    • The abstract and/or introduction should clearly state the claims made, including the contributions made in the paper and important assumptions and limitations. A No or NA answer to this question will not be perceived well by the reviewers.
    • The claims made should match theoretical and experimental results, and reflect how much the results can be expected to generalize to other settings.
    • It is fine to include aspirational goals as motivation as long as it is clear that these goals are not attained by the paper.
  6. 2.
    Limitations
  7. Question: Does the paper discuss the limitations of the work performed by the authors?
  8. Answer: [Yes]
  9. Justification: We discuss limitations of our approach and why they exist in the discussion section 6.
  10. Guidelines:
    • The answer NA means that the paper has no limitation while the answer No means that the paper has limitations, but those are not discussed in the paper.
    • The authors are encouraged to create a separate "Limitations" section in their paper.
    • The paper should point out any strong assumptions and how robust the results are to violations of these assumptions (e.g., independence assumptions, noiseless settings, model well-specification, asymptotic approximations only holding locally). The authors should reflect on how these assumptions might be violated in practice and what the implications would be.
    • The authors should reflect on the scope of the claims made, e.g., if the approach was only tested on a few datasets or with a few runs. In general, empirical results often depend on implicit assumptions, which should be articulated.
    • The authors should reflect on the factors that influence the performance of the approach. For example, a facial recognition algorithm may perform poorly when image resolution is low or images are taken in low lighting. Or a speech-to-text system might not be used reliably to provide closed captions for online lectures because it fails to handle technical jargon.
    • The authors should discuss the computational efficiency of the proposed algorithms and how they scale with dataset size.
    • If applicable, the authors should discuss possible limitations of their approach to address problems of privacy and fairness.
    • While the authors might fear that complete honesty about limitations might be used by reviewers as grounds for rejection, a worse outcome might be that reviewers discover limitations that aren’t acknowledged in the paper. The authors should use their best judgment and recognize that individual actions in favor of transparency play an important role in developing norms that preserve the integrity of the community. Reviewers will be specifically instructed to not penalize honesty concerning limitations.
  11. 3.
    Theory assumptions and proofs
  12. Question: For each theoretical result, does the paper provide the full set of assumptions and a complete (and correct) proof?
  13. Answer: [N/A]
  14. Justification: We do not have theoretical results.
  15. Guidelines:
    • The answer NA means that the paper does not include theoretical results.
    • All the theorems, formulas, and proofs in the paper should be numbered and cross-referenced.
    • All assumptions should be clearly stated or referenced in the statement of any theorems.
    • The proofs can either appear in the main paper or the supplemental material, but if they appear in the supplemental material, the authors are encouraged to provide a short proof sketch to provide intuition.
    • Inversely, any informal proof provided in the core of the paper should be complemented by formal proofs provided in appendix or supplemental material.
    • Theorems and Lemmas that the proof relies upon should be properly referenced.
  16. 4.
    Experimental result reproducibility
  17. Question: Does the paper fully disclose all the information needed to reproduce the main experimental results of the paper to the extent that it affects the main claims and/or conclusions of the paper (regardless of whether the code and data are provided or not)?
  18. Answer: [Yes]
  19. Justification: We provide details to our experimental setup in 4, A, 11. The code is or will be provided.
  20. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • If the paper includes experiments, a No answer to this question will not be perceived well by the reviewers: Making the paper reproducible is important, regardless of whether the code and data are provided or not.
    • If the contribution is a dataset and/or model, the authors should describe the steps taken to make their results reproducible or verifiable.
    • Depending on the contribution, reproducibility can be accomplished in various ways. For example, if the contribution is a novel architecture, describing the architecture fully might suffice, or if the contribution is a specific model and empirical evaluation, it may be necessary to either make it possible for others to replicate the model with the same dataset, or provide access to the model. In general. releasing code and data is often one good way to accomplish this, but reproducibility can also be provided via detailed instructions for how to replicate the results, access to a hosted model (e.g., in the case of a large language model), releasing of a model checkpoint, or other means that are appropriate to the research performed.
    • While NeurIPS does not require releasing code, the conference does require all submissions to provide some reasonable avenue for reproducibility, which may depend on the nature of the contribution. For example
      1. (a)
        If the contribution is primarily a new algorithm, the paper should make it clear how to reproduce that algorithm.
      2. (b)
        If the contribution is primarily a new model architecture, the paper should describe the architecture clearly and fully.
      3. (c)
        If the contribution is a new model (e.g., a large language model), then there should either be a way to access this model for reproducing the results or a way to reproduce the model (e.g., with an open-source dataset or instructions for how to construct the dataset).
      4. (d)
        We recognize that reproducibility may be tricky in some cases, in which case authors are welcome to describe the particular way they provide for reproducibility. In the case of closed-source models, it may be that access to the model is limited in some way (e.g., to registered users), but it should be possible for other researchers to have some path to reproducing or verifying the results.
  21. 5.
    Open access to data and code
  22. Question: Does the paper provide open access to the data and code, with sufficient instructions to faithfully reproduce the main experimental results, as described in supplemental material?
  23. Answer: [Yes]
  24. Justification: All the models, datasets and tasks are openly available. The code is or will be provided.
  25. Guidelines:
    • The answer NA means that paper does not include experiments requiring code.
    • Please see the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • While we encourage the release of code and data, we understand that this might not be possible, so “No” is an acceptable answer. Papers cannot be rejected simply for not including code, unless this is central to the contribution (e.g., for a new open-source benchmark).
    • The instructions should contain the exact command and environment needed to run to reproduce the results. See the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • The authors should provide instructions on data access and preparation, including how to access the raw data, preprocessed data, intermediate data, and generated data, etc.
    • The authors should provide scripts to reproduce all experimental results for the new proposed method and baselines. If only a subset of experiments are reproducible, they should state which ones are omitted from the script and why.
    • At submission time, to preserve anonymity, the authors should release anonymized versions (if applicable).
    • Providing as much information as possible in supplemental material (appended to the paper) is recommended, but including URLs to data and code is permitted.
  26. 6.
    Experimental setting/details
  27. Question: Does the paper specify all the training and test details (e.g., data splits, hyperparameters, how they were chosen, type of optimizer, etc.) necessary to understand the results?
  28. Answer: [Yes]
  29. Justification: We provide details to our experimental setup in 4. Hyperparameters are provided in 11.
  30. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The experimental setting should be presented in the core of the paper to a level of detail that is necessary to appreciate the results and make sense of them.
    • The full details can be provided either with the code, in appendix, or as supplemental material.
  31. 7.
    Experiment statistical significance
  32. Question: Does the paper report error bars suitably and correctly defined or other appropriate information about the statistical significance of the experiments?
  33. Answer: [Yes]
  34. Justification: We provide a study on the robustness of our approach in C.
  35. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The authors should answer "Yes" if the results are accompanied by error bars, confidence intervals, or statistical significance tests, at least for the experiments that support the main claims of the paper.
    • The factors of variability that the error bars are capturing should be clearly stated (for example, train/test split, initialization, random drawing of some parameter, or overall run with given experimental conditions).
    • The method for calculating the error bars should be explained (closed form formula, call to a library function, bootstrap, etc.)
    • The assumptions made should be given (e.g., Normally distributed errors).
    • It should be clear whether the error bar is the standard deviation or the standard error of the mean.
    • It is OK to report 1-sigma error bars, but one should state it. The authors should preferably report a 2-sigma error bar than state that they have a 96% CI, if the hypothesis of Normality of errors is not verified.
    • For asymmetric distributions, the authors should be careful not to show in tables or figures symmetric error bars that would yield results that are out of range (e.g. negative error rates).
    • If error bars are reported in tables or plots, The authors should explain in the text how they were calculated and reference the corresponding figures or tables in the text.
  36. 8.
    Experiments compute resources
  37. Question: For each experiment, does the paper provide sufficient information on the computer resources (type of compute workers, memory, time of execution) needed to reproduce the experiments?
  38. Answer: [Yes]
  39. Justification: We have done an ablation study on the runtime with specified, fixed hardware in 8.
  40. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The paper should indicate the type of compute workers CPU or GPU, internal cluster, or cloud provider, including relevant memory and storage.
    • The paper should provide the amount of compute required for each of the individual experimental runs as well as estimate the total compute.
    • The paper should disclose whether the full research project required more compute than the experiments reported in the paper (e.g., preliminary or failed experiments that didn’t make it into the paper).
  41. 9.
    Code of ethics
  42. Question: Does the research conducted in the paper conform, in every respect, with the NeurIPS Code of Ethics https://neurips.cc/public/EthicsGuidelines?
  43. Answer: [Yes]
  44. Justification: We have conformed to the code of ethics in all aspects.
  45. Guidelines:
    • The answer NA means that the authors have not reviewed the NeurIPS Code of Ethics.
    • If the authors answer No, they should explain the special circumstances that require a deviation from the Code of Ethics.
    • The authors should make sure to preserve anonymity (e.g., if there is a special consideration due to laws or regulations in their jurisdiction).
  46. 10.
    Broader impacts
  47. Question: Does the paper discuss both potential positive societal impacts and negative societal impacts of the work performed?
  48. Answer: [N/A]
  49. Justification: This work advances a subfield in Machine Learning. There are no direct societal impacts caused.
  50. Guidelines:
    • The answer NA means that there is no societal impact of the work performed.
    • If the authors answer NA or No, they should explain why their work has no societal impact or why the paper does not address societal impact.
    • Examples of negative societal impacts include potential malicious or unintended uses (e.g., disinformation, generating fake profiles, surveillance), fairness considerations (e.g., deployment of technologies that could make decisions that unfairly impact specific groups), privacy considerations, and security considerations.
    • The conference expects that many papers will be foundational research and not tied to particular applications, let alone deployments. However, if there is a direct path to any negative applications, the authors should point it out. For example, it is legitimate to point out that an improvement in the quality of generative models could be used to generate deepfakes for disinformation. On the other hand, it is not needed to point out that a generic algorithm for optimizing neural networks could enable people to train models that generate Deepfakes faster.
    • The authors should consider possible harms that could arise when the technology is being used as intended and functioning correctly, harms that could arise when the technology is being used as intended but gives incorrect results, and harms following from (intentional or unintentional) misuse of the technology.
    • If there are negative societal impacts, the authors could also discuss possible mitigation strategies (e.g., gated release of models, providing defenses in addition to attacks, mechanisms for monitoring misuse, mechanisms to monitor how a system learns from feedback over time, improving the efficiency and accessibility of ML).
  51. 11.
    Safeguards
  52. Question: Does the paper describe safeguards that have been put in place for responsible release of data or models that have a high risk for misuse (e.g., pretrained language models, image generators, or scraped datasets)?
  53. Answer: [N/A]
  54. Justification: This paper causes no such risk.
  55. Guidelines:
    • The answer NA means that the paper poses no such risks.
    • Released models that have a high risk for misuse or dual-use should be released with necessary safeguards to allow for controlled use of the model, for example by requiring that users adhere to usage guidelines or restrictions to access the model or implementing safety filters.
    • Datasets that have been scraped from the Internet could pose safety risks. The authors should describe how they avoided releasing unsafe images.
    • We recognize that providing effective safeguards is challenging, and many papers do not require this, but we encourage authors to take this into account and make a best faith effort.
  56. 12.
    Licenses for existing assets
  57. Question: Are the creators or original owners of assets (e.g., code, data, models), used in the paper, properly credited and are the license and terms of use explicitly mentioned and properly respected?
  58. Answer: [Yes]
  59. Justification: All models, tasks, datasets and prior pruning techniques have been appropriately cited.
  60. Guidelines:
    • The answer NA means that the paper does not use existing assets.
    • The authors should citep the original paper that produced the code package or dataset.
    • The authors should state which version of the asset is used and, if possible, include a URL.
    • The name of the license (e.g., CC-BY 4.0) should be included for each asset.
    • For scraped data from a particular source (e.g., website), the copyright and terms of service of that source should be provided.
    • If assets are released, the license, copyright information, and terms of use in the package should be provided. For popular datasets, paperswithcode.com/datasets has curated licenses for some datasets. Their licensing guide can help determine the license of a dataset.
    • For existing datasets that are re-packaged, both the original license and the license of the derived asset (if it has changed) should be provided.
    • If this information is not available online, the authors are encouraged to reach out to the asset’s creators.
  61. 13.
    New assets
  62. Question: Are new assets introduced in the paper well documented and is the documentation provided alongside the assets?
  63. Answer: [N/A]
  64. Justification: The paper does not release new assets.
  65. Guidelines:
    • The answer NA means that the paper does not release new assets.
    • Researchers should communicate the details of the dataset/code/model as part of their submissions via structured templates. This includes details about training, license, limitations, etc.
    • The paper should discuss whether and how consent was obtained from people whose asset is used.
    • At submission time, remember to anonymize your assets (if applicable). You can either create an anonymized URL or include an anonymized zip file.
  66. 14.
    Crowdsourcing and research with human subjects
  67. Question: For crowdsourcing experiments and research with human subjects, does the paper include the full text of instructions given to participants and screenshots, if applicable, as well as details about compensation (if any)?
  68. Answer: [N/A]
  69. Justification: This work does not involve crowdsourcing nor research with human subjects.
  70. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Including this information in the supplemental material is fine, but if the main contribution of the paper involves human subjects, then as much detail as possible should be included in the main paper.
    • According to the NeurIPS Code of Ethics, workers involved in data collection, curation, or other labor should be paid at least the minimum wage in the country of the data collector.
  71. 15.
    Institutional review board (IRB) approvals or equivalent for research with human subjects
  72. Question: Does the paper describe potential risks incurred by study participants, whether such risks were disclosed to the subjects, and whether Institutional Review Board (IRB) approvals (or an equivalent approval/review based on the requirements of your country or institution) were obtained?
  73. Answer: [N/A]
  74. Justification: This work does not involve crowdsourcing nor research with human subjects.
  75. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Depending on the country in which research is conducted, IRB approval (or equivalent) may be required for any human subjects research. If you obtained IRB approval, you should clearly state this in the paper.
    • We recognize that the procedures for this may vary significantly between institutions and locations, and we expect authors to adhere to the NeurIPS Code of Ethics and the guidelines for their institution.
    • For initial submissions, do not include any information that would break anonymity (if applicable), such as the institution conducting the review.
  76. 16.
    Declaration of LLM usage
  77. Question: Does the paper describe the usage of LLMs if it is an important, original, or non-standard component of the core methods in this research? Note that if the LLM is used only for writing, editing, or formatting purposes and does not impact the core methodology, scientific rigorousness, or originality of the research, declaration is not required.
  78. Answer: [N/A]
  79. Justification: The core method development did not involve LLMs.
  80. Guidelines:
    • The answer NA means that the core method development in this research does not involve LLMs as any important, original, or non-standard components.
    • Please refer to our LLM policy (https://neurips.cc/Conferences/2025/LLM) for what should or should not be described.