### v2

# TRIM: Achieving Extreme Sparsity with Targeted Row-wise Iterative Metric-driven Pruning

Abstract
Large Language Models (LLMs) present significant computational and memory challenges due to their extensive size, making pruning essential for their efficient deployment. Existing one-shot pruning methods often apply uniform sparsity constraints across layers or within each layer, resulting in suboptimal performance at high sparsity ratios. This work introduces TRIM (Targeted Row-wise Iterative Metric-driven pruning), a novel approach that applies varying sparsity ratios to individual output dimensions (rows) within each layer. TRIM employs an iterative adjustment process guided by quality metrics to optimize dimension-wise sparsity allocation, ensuring a more equal quality retention across output dimensions to preserve critical information. TRIM can be seamlessly integrated with existing layer-wise pruning strategies. Our evaluations on perplexity and zero-shot tasks across diverse LLM families (Qwen2.5, LLaMA-2, and OPT) and sparsity levels demonstrate that TRIM achieves new state-of-the-art results and enhances stability. For instance, at 80% sparsity, TRIM reduces perplexity by 48% for Qwen2.5-14B and over 90% for OPT-13B compared to baseline methods. We conclude that fine-grained, dimension-wise sparsity adaptation is crucial for pushing the limits of extreme LLM compression. Code available at: github.com/flobk/TRIM
TRIM: Achieving Extreme Sparsity with Targeted Row-wise Iterative Metric-driven Pruning
Florentin Beck1, William Rudman2, Carsten Eickhoff1, 1 University of Tübingen, School of Medicine 2 University of Texas at Austin, Department of Linguistics florentin.beck@student.uni-tuebingen.de
1 Introduction
Large language models (LLMs) have gained prominence due to their generalized problem-solving abilities (Brown et al., 2020; OpenAI et al., 2024; Zhao et al., 2025). A significant driver of this success has been the exponential scaling of model parameters (Kaplan et al., 2020; Hoffmann et al., 2022). Increasing the number of parameters has improved LLM performance and enabled emergent behaviors (Wei et al., 2022), such as in-context learning (Brown et al., 2020) and reasoning (Yao et al., 2023; Wei et al., 2023). However, this growth in model size has lead to substantial memory and computational demands, which pose considerable deployment challenges (Zhou et al., 2024).
As LLM parameters scale exponentially, pruning techniques (Mozer and Smolensky, 1988; Le Cun et al., 1989; Han et al., 2015) have become increasingly important to enable efficient inference in resource-constrained settings. The goal of pruning is to reduce the number of model weights while preserving performance on downstream tasks, which helps alleviate GPU memory requirements.
Several pruning techniques have been developed to make LLMs more efficient. One major class induces sparsity during training or with fine-tuning, dynamically removing less important weights (Han et al., 2015; Evci et al., 2021; Louizos et al., 2018). While effective, these methods often require modifications to the training process and can be computationally expensive, particularly for models with billions of parameters, where maintaining optimizer states and gradients becomes a significant bottleneck (Molchanov et al., 2019; Liu et al., 2019; Hoefler et al., 2021; Gale et al., 2019).
As a result, one-shot pruning methods that reduce LLM parameters without retraining to compute weight masks have gained prominence (Han et al., 2015; Sun et al., 2024; Zhang et al., 2024a; Das et al., 2024). A challenge with current approaches is that they often impose predefined structural constraints, such as uniform sparsity across all network layers or outputs. This imposed structure is arbitrary and can be detrimental to pruning performance, particularly as LLMs exhibit unique weight and activation characteristics, such as prominent outlier features and highly skewed activation distributions (Kovaleva et al., 2021; Dettmers et al., 2022; Rudman et al., 2024).
Figure 1: Illustrating non-uniform, dimension-wise sparsity. On the left is Wanda, which applies the layer sparsity ratio $\mathbf{T}$ uniformly to all output dimension (rows) of the weight matrix $\mathbf{W}$. TRIM iteratively defines sparsity ratios for individual dimensions in a non-uniform way. This targeted distribution of the available sparsity budget improves local (and global) pruning quality.
Layer-wise sparsity adaptation methods aim to mitigate this by assigning distinct sparsity ratios to each layer of the model, determined through algorithmic search (Li et al., 2024; Xu et al., 2025), heuristic-based approaches (Yin et al., 2024; Sun et al., 2025), or theoretically grounded strategies (Lu et al., 2024). While this offers more flexibility than uniform sparsity, these methods still apply a consistent sparsity target within each layer and can struggle under extreme sparsity levels, such as removing more than 70% of model parameters.
In this work, we propose TRIM: (Targeted Row-wise Iterative Metric-driven pruning). TRIM builds on the principles of one-shot pruning algorithms like Wanda (Sun et al., 2024), which compare and remove weights at each layer iteratively. Current methods typically apply a uniform sparsity constraint across all output dimensions (per-output), however, TRIM introduces a more granular approach. TRIM calculates unique sparsity ratios for individual output dimensions (corresponding to a row in the layer’s weight matrix). This fine-grained allocation is guided by the objective of reducing variance in quality retention across different outputs during the pruning process. We demonstrate that TRIM achieves state-of-the-art pruning performance at extreme sparsity levels (e.g., upwards of 80% of parameters removed).
Because TRIM adapts sparsity at the dimension-wise level, it is inherently compatible with and can augment existing layer-wise allocation strategies. In this paper, we use TRIM to improve two layer-wise pruning methods: OWL and AlphaPruning.
TRIM readily integrates into importance-score-based pruning algorithms, providing a principled extension of such methods beyond any single metric (e.g., not limited to Wanda). In doing so, TRIM preserves computational efficiency and broadens the applicability of per-output pruning approaches. The contributions of our work are as follows:
  1. 1.
    We introduce TRIM, the first algorithm to enhance pruning granularity by calculating sparsity ratios on a per-dimension basis.
  2. 2.
    Using TRIM, we preserve performance at lower sparsity ratios and achieve state-of-the-art pruning results at extreme levels of sparsity reducing the perplexity of OPT-13B at 80% from 6461 with Wanda-based OWL to 324 with TRIM.
  3. 3.
    We conduct a thorough empirical analysis as to why TRIM leads to superior pruning results and show that individual output dimensions differ in their 1) sensitivity to pruning and 2) criticality for down-stream performance.
2 Related Work
A large body of work has explored pruning techniques that rely on gradient information or require retraining to recover performance. Gradient-based methods such as SNIP (Lee et al., 2019), GraSP (Wang et al., 2020), and SynFlow (Tanaka et al., 2020) prune weights based on sensitivity estimates derived from gradients at initialization. Iterative methods like OBD (Le Cun et al., 1989), OBS (Hassibi et al., 1993), and movement pruning (Sanh et al., 2020) leverage gradients or Hessians during or after training to inform pruning decisions, typically requiring retraining or fine-tuning to regain lost performance (Molchanov et al., 2019; Liu et al., 2019; Frankle and Carbin, 2019).
While these methods achieve strong results, the computational burden of retraining models with billions of parameters has driven the development of efficient one-shot pruning approaches that eliminate the need for retraining. Zhang et al. (2024b); van der Ouderaa et al. (2024)
[…]
Methods such as Outlier Weighed Layerwise (OWL) sparsity extend Wanda by introducing layer-wise sparsity ratios (Yin et al., 2024). OWL assigns unique sparsity ratios to individual layers of a model while maintaining a global sparsity target. By assigning unique sparsity ratios to each layer, OWL avoids pruning layers that contain many outlier positions. An outlier position exists when its Wanda score, $A_{ij}$, is greater than a predefined multiple, $M$, of the mean Wanda score of that layer, $\bar{A}$. Namely, if $A_{ij}>M\cdot\bar{A}$. AlphaPruning (Lu et al., 2024) offers an alternative to OWL by assigning layer-wise sparsity ratios based on the heavy-tailed shape of spectral densities in layer weight matrices. It has its roots in heavy-tailed self-regularization theory (Martin and Mahoney, 2019a, b), making it a more principled approach than OWL.
Current pruning methods, such as Wanda and its derivations OWL, and AlphaPruning, apply a uniform sparsity ratio to all output dimensions within a given layer. We demonstrate that this uniform allocation is not optimal when pruning at high levels of sparsity because different output dimensions have varying sensitivities to pruning. To address this, we propose a more granular approach by relaxing the constraint of uniform sparsity per dimension. Instead, TRIM assigns varying sparsity ratios to each output dimension $W_{i,:}$ of a weight matrix. Importantly, while the sparsity ratio can differ per output dimension, the original mechanism for selecting which weights to prune within that dimension (up to its assigned ratio) remains the same as the existing algorithm TRIM adapts.
Algorithm 1 Iterative Dimension-Wise Sparsity Adjustment
[…]
3 TRIM
Table 1: Perplexity results on WikiText validation set for various models pruned at 70% and 80% sparsity. The $\Delta$ column shows the average perplexity reduction across the tested models achieved by adding TRIM. Lower is better.
Method OPT LLaMA-2 Qwen2.5 6.7B 13B 7B 13B 3B 7B 14B 32B 72B $\Delta$ 0% - 10.86 10.13 5.47 4.88 8.03 6.85 5.29 5.018 3.851 - 70% Alpha 63.53 820.42 31.32 15.19 164.39 68.93 61.59 15.70 11.66 -25.3% +TRIM 30.80 35.16 29.00 14.70 128.44 62.22 46.43 13.94 11.41 OWL 40.30 28.92 21.20 16.02 83.98 44.60 36.46 13.06 11.25 -4.6% +TRIM 38.21 28.21 20.10 14.39 83.39 43.97 33.21 12.28 11.06 80% Alpha 5357.60 4726.09 1558.41 143.19 9872.21 728.82 1040.91 469.52 70.26 -41.6% +TRIM 656.26 557.30 1527.01 108.70 1685.51 1037.11 435.84 133.94 68.64 OWL 8509.51 6461.43 406.20 225.04 768.69 245.32 348.48 73.44 61.59 -34.9% +TRIM 492.47 324.14 437.88 154.83 504.89 233.80 180.67 66.95 58.15
Figure 1 illustrates the core concept of TRIM. Let $W\in\mathbb{R}^{D\times N}$ be the weight matrix of a layer, where $D$ is the number of output dimensions (rows) and $N$ is the number of input dimensions (columns). Each row $W_{i,:}$ corresponds to the weights associated with the $i$-th output dimension. We define a dimension-wise sparsity vector $S=[S_{1},S_{2},\dots,S_{D}]$, where each element $S_{i}\in[0,1]$ specifies the target sparsity ratio for the $i$-th output dimension $W_{i,:}$. This means that for each row $W_{i,:}$, $S_{i}\cdot N$ of its weights will be pruned. While the individual sparsity ratios $S_{i}$ can vary, they must collectively satisfy the target sparsity $T$ for the layer. This is enforced by ensuring their average equals $T$:
$\frac{1}{D}\sum_{i=1}^{D}S_{i}=T$ (1)
Setting $S_{i}=T$ for all $i\in\{1,\dots,D\}$ reverts to the standard pruning approach where sparsity is uniform across output dimensions.
The core challenge is to determine an effective allocation strategy for the dimension-wise sparsity vector, $S$, that optimizes performance while adhering to Equation 1. Below, we detail our method for finding $S$, given a layer’s weight matrix, $W$, and $L$ sample input activations $X\in\mathbb{R}^{N\times L}$.
Calculating Dimension-Wise Sparsity Vectors.
To find a suitable dimension-wise sparsity vector $S$, we employ an iterative adjustment process detailed in Algorithm 1. The process begins by computing the unpruned output $Y\leftarrow WX$ (Line 4), with $Y\in\mathbb{R}^{D\times L}$ and initializing $S$ uniformly, so that $S_{i}=T$ for all output dimensions $i=1,\dots,D$ (Line 5). This initial $S$ also serves as the first candidate for the best sparsity vector, $S_{best}$. The algorithm then iterates $K$ times (Lines 8-20). In each iteration $k$:
  1. 1.
    The weight matrix $W$ is pruned using the current sparsity vector to obtain $W_{pruned}$ (Line 9). This means for each row $W_{i,:}$, $S_{i}\cdot N$ weights are removed based on the chosen pruning criterion (e.g., lowest Wanda scores).
  2. 2.
    The pruned output $\hat{Y}\leftarrow W_{pruned}X$ is computed (Line 10).
  3. 3.
[…]
  4. 4.
    The quality $c_{i}\leftarrow\text{QmetricDimwise}(Y_{i,:},\hat{Y}_{i,:})$ for each individual output dimension is calculated (Line 15).
  5. 5.
    The sparsity targets $S_{i}$ are adjusted based on these per-dimension quality scores. First, the scores $c_{i}$ are normalized to a [0, 1] range (Line 16). An update term $\delta_{i}\leftarrow\alpha c^{\prime}_{i}$ is calculated using a learning rate $\alpha$ (Line 17). This mechanism aims to increase the sparsity target $S_{i}$ for dimensions exhibiting higher quality (less degradation, higher $c^{\prime}_{i}$) and decrease it for dimensions with lower quality, thereby pruning less sensitive dimensions more aggressively.
  6. 6.
    Finally, the adjusted sparsity vector is re-centered to ensure its average value remains equal to the overall target layer sparsity $T$ (Line 18) in order to satisfy Equation 1.
After $K$ iterations, the algorithm returns $S_{best}$ (Line 21), which represents the dimension-wise sparsity allocation that yields the highest overall quality $q_{best}$ during the iterative search. $S_{best}$ is then used to perform the final pruning of the layer. In Appendix G, we explore the impact of #$k$ on the pruning quality of TRIM.
To determine the optimal sparsity distribution, we evaluate positive learning rates $\alpha$, beginning with small values, applying Algorithm 1, and recording $q_{best}$. We increase $\alpha$ and repeat the process, stopping when a larger $\alpha$ no longer improves $q_{best}$. However, if a positive $\alpha$ does not yield a quality $q_{best}$ superior to the uniform sparsity baseline, we use a negative learning rate. This allows for better generalization across different model activation spaces. In Section 5, we briefly explore when a negative learning rate occurs in models.
We use cosine similarity as our quality metric and refer the reader to Appendix A for the evaluation of other metrics. Additionally, Appendix B presents our analysis of the computational overhead of this iterative procedure. Our findings show that calculating the sparsity vectors for all layers of the Qwen2.5-14B model adds only a small delay, taking between 27.76 and 59.88 seconds based on the number of calibration samples used.
Figure 2: Perplexity progression from 70% to 80% sparsity. TRIM extends the usable sparsity range.
4 Experiments
Models and Evaluation.
We evaluate TRIM on Qwen2.5 3B/7B/14B/32B/72B (Qwen et al., 2025), LLaMA-2 7B/13B (Touvron et al., 2023), and OPT 7B/13B (Zhang et al., 2022). Our evaluation adheres to established methodologies for pruning LLMs (Sun et al., 2024; Frantar and Alistarh, 2023). We evaluate language modeling proficiency by calculating perplexity on the reserved WikiText validation set (Merity et al., 2016) and assess zero-shot performance across a range of down-stream tasks. We evaluate on BoolQ (Clark et al., 2019), RTE (Wang et al., 2019), HellaSwag (Zellers et al., 2019), WinoGrande (Sakaguchi et al., 2019), ARC Easy and Challenge (Clark et al., 2018), and OpenbookQA (Mihaylov et al., 2018), using the framework by Gao et al. (Mihaylov et al., 2018). Additionally, we evaluate the generalizability of the WikiText results on C4 and Pile in Appendix F.
Baselines.
We evaluate Wanda (Sun et al., 2024) based OWL (Yin et al., 2024) and AlphaPruning (Lu et al., 2024) with and without TRIM. In Appendix C, we evaluate Wanda with uniform layerwise sparsities. For all runs, we use the same number of calibration samples, each with a sequence length of 2048 tokens, randomly selected from the first shard of the C4 dataset (Raffel et al., 2023). Hyperparameters can be found in Appendix H.
Table 2: Average zero-shot accuracies (%) for models pruned at 70% and 80% sparsity. $\Delta$ shows the average percentage increase achieved by adding TRIM. Higher is better.
Method OPT LLaMA-2 Qwen2.5 6.7B 13B 7B 13B 3B 7B 14B 32B 72B $\Delta$ 0% - 51.52 52.60 59.73 63.02 61.13 65.82 68.12 68.09 70.13 - 70% Alpha 38.42 39.35 42.01 47.39 35.43 39.65 39.92 52.98 60.80 +1.14% +TRIM 42.27 42.27 41.60 47.92 36.75 39.75 40.73 53.15 61.76 OWL 40.67 43.60 43.30 49.01 36.53 39.22 40.92 54.56 61.30 +0.42% +TRIM 41.16 43.12 43.36 49.54 36.65 39.82 41.68 55.04 62.47 80% Alpha 32.70 35.13 31.94 36.71 32.77 32.74 35.32 36.06 40.32 +0.65% +TRIM 35.67 36.56 31.99 36.73 33.16 32.98 34.40 37.44 40.59 OWL 32.61 35.22 32.60 36.23 35.42 32.31 33.70 37.35 41.19 +0.46% +TRIM 36.30 35.47 32.79 36.87 34.82 32.64 33.32 37.48 41.02
4.1 Main results
We evaluate OWL and AlphaPruning with and without TRIM at high sparsity ratios (70% and 80%), and additionally evaluate OWL with and without TRIM at 50% and 60%. For all methods, we only prune linear layers, excluding the initial embedding layer and the final classification head.
Language Modeling.
Table 1 shows perplexity results for all models used in this study. Integrating TRIM with OWL and AlphaPruning consistently results in lower perplexity values than these baselines without TRIM. At 70% and 80% sparsity respectively, TRIM reduces perplexity for LLaMA-2-13B by 10% and 31%, and for Qwen2.5-14B by 9% and 48%. The effectiveness of TRIM is especially pronounced for OPT-13B at 80% sparsity, where TRIM decreases OWL pruning perplexity from 6461.43 to 324.14.
TRIM improves the robustness and reliability of pruning results, correcting counterintuitive behaviors seen with baseline methods. At 80% sparsity, Qwen2.5-14B shows higher perplexity than its smaller 7B counterpart, and OPT models suffer from unusually poor performance. TRIM resolves both issues, restoring expected scaling and stable performance. These consistent and robust improvements across all model families, sizes, and sparsity levels underscore the critical role of effective dimension-wise sparsity allocation and highlight TRIM’s efficacy in achieving it.
Table 3: Perplexity results on WikiText validation set for various models pruned at 50% and 60% sparsity. The $\Delta$ column shows the average perplexity reduction across the tested models achieved by adding TRIM. Lower is better.
Method OPT LLaMA-2 Qwen2.5 6.7B 13B 7B 13B 3B 7B 14B 32B 72B $\Delta$ 50% OWL 12.39 12.10 7.24 6.09 12.04 8.80 8.16 6.31 5.29 -0.8% +TRIM 12.39 12.08 7.15 5.99 11.94 8.79 8.12 6.32 5.27 60% OWL 16.12 16.39 9.65 7.82 24.91 9.65 11.67 7.64 6.70 -2.3% +TRIM 16.04 16.38 9.36 7.49 23.53 9.35 11.36 7.58 6.63
Zero-Shot Downstream Tasks.
The benefits of TRIM extend beyond perplexity improvements to generalization on downstream tasks. As evidenced by the average zero-shot accuracies presented in Table 2, integrating TRIM with layer-wise pruning strategies yields superior performance compared to baseline methods without TRIM. This performance advantage holds across a wide range of models and sparsity levels. The complete result tables can be found in Appendices 16 and 17.
4.2 Lower Sparsity Levels.
At moderate sparsity (50-60%), TRIM matches or improves WikiText validation perplexity relative to the OWL baseline across all tested models. These results (Table 3) indicate that TRIM’s benefits persist below the high-sparsity regime, enhancing or preserving quality also at lower pruning ratios.
4.3 Generalization To Other Pruning Metrics.
Although we primarily evaluate TRIM with Wanda (Sun et al., 2024), its dimension-wise sparsity and per-output optimization apply to any scoring rule that assigns a value to each weight. For clarity, the three metrics we test are: (i) Magnitude, which uses absolute weights $|\mathbf{W}_{ij}|$ Han et al. (2015); (ii) SparseGPT, which scores weights by the Hessian-aware ratio $\lvert\mathbf{W}\rvert^{2}/\mathrm{diag}(\mathbf{H}^{-1})]_{ij}$ with $\mathbf{H}\!\approx\!\mathbf{X}^{\top}\mathbf{X}$ from layer inputs $\mathbf{X}$ (Frantar and Alistarh, 2023); (iii) GBLM which extends Wanda by multiplying weight magnitude with an activation- and gradient-based term, $|\mathbf{W}_{ij}|\!\cdot\!(\lVert\mathbf{X}_{j}\rVert+\alpha\,\mathbf{G}_{ij})$ (i.e., Wanda + gradient score) (Das et al., 2024).
SparseGPT compares candidates over blocks of 128 inputs (columns), Magnitude compares layerwise, whereas Wanda/GBLM compare within a single output (row). We omit SparseGPT’s post-pruning weight updates to isolate the effect of pruning configuration. Table 4 shows that TRIM’s per-output configuration yields the best perplexity across all three metrics, indicating that TRIM’s gains generalize well beyond the Wanda metric.
Table 4: WikiText perplexity of LLaMA-2-13B at 70% with OWL layer ratios. TRIM improves upon the per-output configuration and generalizes to other pruning metrics. The default comparison group is underlined.
Method (layer) (128, input) (1, output) / +TRIM Magnitude 59.86 49.25 23.48 / 19.94 SparseGPT 19.75 19.23 15.19 / 14.11 GBLM 18.53 22.80 16.04 / 14.15
(a) Gini histogram.
(b) Cosine similarity vs Sparsity.
(c) LR sign vs Gini (outlier concentration).
Figure 3: Gallery of pruning diagnostics. (a) Gini histogram: Gini coefficients of the Wanda pruning metric across output dimensions. Higher Gini $\Rightarrow$ signal concentrated in fewer weights. (b) Cosine similarity vs Sparsity: Dimension-wise cosine similarities at increasing sparsity, showing heterogeneous pruning sensitivity across dimensions. (c) LR sign vs outlier Gini: LLaMA-2-13B (higher Gini; more concentrated outliers) uses more negative effective learning rates than Qwen2.5-14B (lower Gini; more uniform outliers). Both plots show the gate-projection.
5 Analysis
In this section, we highlight three key empirical observations that explain TRIM’s effectiveness: 1) output dimensions exhibit varying sensitivity to pruning; 2) post-pruning quality degrades as sparsity levels increase; and 3) some output dimensions are critical to model performance, while others contribute little to model performance.
1. Output dimensions exhibit different sensitivity to pruning.
[…]
2. Quality degradation accelerates with increasing sparsity.
Pruning can remove redundant weights with minimal impact at low sparsity levels. However, as sparsity rises, pruning impacts weights with essential activation signals, leading to a faster decay in performance. This non-linear decline is evident in the average trend of quality degradation on the dimension-level (Figure 3(b)) and in downstream metrics like the exponential increase in perplexity at high sparsity rates (Figure 2).
The heterogeneity of importance concentration translates directly into differing pruning sensitivities across output dimensions. To demonstrate this, we increase the uniform sparsity ratio applied to a layer and measure the resulting quality degradation for each individual output dimension. We quantify the quality using the cosine similarity $c_{i}=\text{CosineSimilarity}(Y_{i,:},\hat{Y}_{i,:})$ between the original output $Y_{i,:}$ and the pruned output $\hat{Y}_{i,:}$ for that dimension, given a set of sample inputs $X$. Figure 3(b) plots these cosine similarities $c_{i}$ against increasing levels of sparsity $T$. The rate at which quality degrades differs significantly across dimensions. Some dimensions maintain high similarity even at high sparsity levels (resilient to pruning), while others experience a rapid drop even at lower sparsity (sensitive to pruning).
Table 5: Wikitext validation perplexity for Qwen2.5-3B. Exactly one dimension gets removed.
Method Perplexity Baseline 8.03 $D$ w. min norm 8.19 $D$ w. max norm 273.10 Random $D$ 8.18
3. Output dimensions differ in their impact on overall performance.
Beyond the differences in their sensitivity to pruning, output dimensions vary significantly in their importance for model performance. To illustrate this, we fully eliminate exactly one output dimension per layer in Qwen2.5-3B, resulting in 0.03% total sparsity (Table 5). Removing the dimension with the smallest L2-norm only increases perplexity by 0.16 from the unpruned baseline, similar to choosing a random dimension (+0.15). However, pruning the dimension with the highest norm increases perplexity to 273.10.
The importance of a dimension is also linked to the concentration of outlier positions present in it. Recall that $A_{ij}$ is an outlier position if $A_{ij}>M\cdot\bar{A}$. We take the top 10% of dimensions with the largest number of outliers (i.e., "outlier-dense" dimensions) and prune them with 90% sparsity, resulting in 9% total model sparsity. Note that pruning this way retains all outlier positions since they are always the last weights to get pruned. Table 6 shows that pruning outlier-dense dimensions significantly impacts performance compared to randomly pruning an equivalent number of dimensions at the same sparsity. This result highlights two key points.
Table 6: Wikitext validation perplexity. 10% of dimensions $D$ are heavily pruned (90%). Outlier-dense dimensions are important for performance.
Model Baseline Random Outlier-dense $D$ Qwen2.5-7B 6.85 9.17 26.39 Qwen2.5-14B 5.29 7.70 26.16 LLaMA2-7B 5.47 7.27 255.06 LLaMA2-13B 4.88 6.23 74.27
First, outlier-dense dimensions are important for model performance. Second, although outlier positions in dense dimensions appear to capture most of the signal, non-outlier weights retain substantial information, limiting aggressive pruning. This effect varies across model families. For example, LLaMA models exhibit a more pronounced degradation than Qwen, indicating that outlier-dense dimensions are more important for some models than for others. Together, these findings highlight that output dimensions are not uniformly important and that simple heuristics, such as L2-Norm or outlier characteristics, can identify dimensions critical for preserving model capabilities.
6 Discussion
Targeted sparsity allocation has proven successful for enhancing the performance of pruned LLMs, but has so far only been attempted at the layer-wise level (Yin et al., 2024; Lu et al., 2024; Sun et al., 2025; Xu et al., 2025). This work extends this paradigm to individual output dimensions within each layer. Our findings show that dimension-wise sparsity is necessary, when pruning at high sparsity levels. Observation 5 demonstrates that post-pruning quality tends to degrade at an accelerated rate as sparsity increases. As a result, using coarse pruning strategies becomes increasingly detrimental. The core premise of TRIM is that significant heterogeneity exists at the dimension level (Observations 5 and 5). However, moving to such a fine-grained level of control introduces challenges: the search space for optimal sparsity configurations becomes substantially larger, and the feedback signals (e.g., quality metrics per dimension) might become less reliable or more noisy compared to layer-level aggregates.
TRIM addresses this complexity with two primary ideas. First, reducing the variance in post-pruning quality degradation across dimensions is beneficial for overall layer performance. This is pursued via a positive learning rate in its iterative refinement process, aiming to prevent any single dimension from experiencing significant quality loss. Second, TRIM recognizes that simple quality equalization is not universally optimal. Some output dimensions possess both high initial post-pruning quality and are critical to model function (e.g., outlier-dense dimensions; Table 6). Forcing these dimensions to match the potentially lower quality of less important, more sensitive ones through aggressive variance reduction would be counterproductive. This leads to TRIM’s adaptive mechanism, which involves the monitoring of layer quality and the learning rate adaptation. When beneficial, TRIM can use a negative learning rate, effectively increasing quality variance. LLaMA models, for instance, often exhibit outlier-dense dimensions (reflected in high Gini coefficients for outlier concentration, Figure 3(c)) that are important for model performance (Table 6). As a result, TRIM frequently selects a negative learning rate for these models (approximately 40% of the time, compared to 10% for other models).
7 Conclusion
In this paper, we introduce TRIM (Targeted Row-wise Iterative Metric-driven pruning), a novel approach that, extends targeted sparsity allocation not only across layers but within each layer by introducing the concept of dimension-wise sparsity. TRIM enables stable performance in challenging pruning scenarios by iteratively refining a sparsity ratio for each output. Using TRIM, we reduce perplexity by 48% for Qwen2.5-14B and over 90% for OPT-13B compared to existing state-of-the-art methods when pruning at 80% sparsity. Beyond these gains, TRIM is a lightweight, future-proof extension to importance-score pruning: it decouples scoring from allocation and accepts any importance metric as input. This plug-and-play design layers dimension-wise budgets on top of existing scorers, improving high-sparsity stability with negligible overhead, and providing a clear upgrade path as new scoring rules emerge.
Limitations
Layers with concentrated outliers benefit from negative learning rates $\alpha$, whereas more uniform layers prefer positive $\alpha$ (Fig. 3(c), Tab. 6). These findings present an inherent tension between the general benefit of reducing quality variance for overall stability and the need to preserve and potentially increase variance in highly critical dimensions. TRIM’s flexible learning rate and best-state selection address this tension, but do not eliminate it; exploring alternative update rules or explicit variance-stability trade-offs could be promising for future research.
Our allocation is unstructured within each row, and it is this flexibility that allows TRIM to outperform previous methods. However, this freedom also means it does not directly translate to fixed-sparsity, hardware-accelerated patterns such as n:m.
Appendix A Quality Metric
TRIM utilizes two distinct quality evaluations to guide the pruning process. The first, layer-wise quality evaluation, assesses the pruning quality for an entire layer. We conduct an ablation study across different sparsity levels, model sizes, and model families, with results presented in Table 7. The metrics considered are Mean Squared Error (MSE), Cosine Flat (which flattens the two-dimensional output vectors before calculating cosine similarity), and Cosine Sample (which calculates cosine similarity along the sample axis and then averages the results). Our findings indicate that both Cosine metrics offer a robust balance between accuracy and reliability across diverse pruning scenarios.
The second evaluation, dimension-wise quality assessment, measures the post-pruning quality for each individual output dimension. For this assessment, we investigate Cosine Similarity, Mean Squared Error (MSE), and Peak Signal-to-Noise Ratio (PSNR). As shown in Table 8, while MSE can outperform Cosine Similarity in specific cases (e.g., for LLaMA models), Cosine Similarity generally demonstrates greater reliability.
These results demonstrate that: a) Cosine Similarity, as the default configuration for both layer-wise and dimension-wise quality assessments, proves to be a reliable choice, b) TRIM’s performance can be further optimized in specific pruning scenarios by performing a hyperparameter sweep to identify the most suitable quality metric, and c) the choice of dimension-wise quality metric is typically more important than the layer-wise metric for overall pruning performance.
Appendix B Runtime Overhead
One-shot pruning methods like Wanda (Sun et al., 2024) are designed for efficient pruning of models with billions of parameters. Any extension to these methods should maintain this computational efficiency. TRIM, while iterative in nature, is designed around that and performs most computations efficiently vectorized on the GPU.
To empirically validate this claim, we prune the Qwen2.5-14B model using one NVIDIA A100-40GB and report the time in Table 9. TRIM adds little overhead compared to the duration of the entire Wanda based pruning function (8% runtime increase for 32 samples, and 1.8% increase for 512 samples).
Table 10: Perplexity results on WikiText validation set for various models pruned at 50% and 60% sparsity. The $\Delta$ column shows the average perplexity reduction across the tested models achieved by adding TRIM. Lower is better.
Method OPT LLaMA-2 Qwen2.5 6.7B 7B 13B 3B 7B 14B $\Delta$ 0% - 10.86 5.47 4.88 8.03 6.85 5.29 - 50% Wanda 11.96 6.92 5.98 11.37 8.57 7.30 -0.94% +TRIM 11.98 6.88 5.96 11.18 8.55 7.29 60% Wanda 15.14 10.77 8.41 21.51 13.75 11.31 -11.42% +TRIM 14.68 10.64 8.26 19.06 13.05 9.98
Table 11: Robustness analysis of TRIM with different random seeds on Qwen2.5-14B and LLaMA-2-7B at 70% sparsity. Wikitext validation perplexity, lower is better.
Model Method Seed 1 Seed 2 Seed 3 Seed 4 Seed 5 Mean Std
Qwen2.5-14B OWL 37.63 37.43 36.33 37.71 36.87 37.20 0.52
Qwen2.5-14B +TRIM 33.64 33.50 33.61 33.16 34.95 33.78 0.61
LLaMA-2-7B OWL 21.53 21.39 21.32 21.58 21.60 21.48 0.11
LLaMA-2-7B +TRIM 20.44 20.37 19.95 20.40 20.73 20.38 0.25
Appendix C TRIM without Layerwise Sparsity Allocation
In the main body of the paper, we evaluate TRIM in conjunction with layerwise sparsity allocation schemes such as OWL Yin et al. (2024) and AlphaPruning Lu et al. (2024). To test if the improvements of TRIM hold up without these schemes, we evaluate several models on the Wanda metric with uniform sparsity ratios (see Table 10). The improvements of TRIM without layerwise allocation schemes are comparable to the improvements with layerwise allocation schemes (compared to Table 3). In fact, since the baseline performance is typically worse without the layerwise schemes, TRIM typically demonstrates larger relative performance gains.
Appendix D Robustness Analysis
To evaluate the robustness, we run TRIM five times with different random seeds and report the results in Table 11. TRIM slightly increases the standard deviation compared to the OWL baseline method.
Appendix E Implementation Details
Sparsity Cutoff.
The limited amount of calibration data can push the sparsity for a dimension into a local minimum, sometimes completely eliminating it. To prevent this overfitting from happening, we limit the sparsity that any dimension can experience to at most 95%. This helps TRIM achieve more stable results.
Input Recalculation.
Additionally, we give the option for input vector recalculation. One-shot pruning methods like Wanda prune block by block, and in each block layer by layer. However, when the Q,K,V-projections are pruned the activation statistic for subsequent layers (Out-projection, or the MLP block) also changes. Therefore, we include an option to recalculate the input vectors used by TRIM. Note that this does not include the Wanda metric itself, which ensures fair comparisons. This recalculation makes pruning quality slightly better (see Table 12) and results in increased pruning time. We recommend enabling this option for achieving the best results, and disabling it during hyperparameter tuning or developing.
Table 12: Effect of input recalculation on perplexity for Qwen2.5-14B at different sparsity levels.
Sparsity Baseline Without Recalc With Recalc 70% 36.46 33.85 33.21 80% 348.48 186.32 180.67
Appendix F Generalizability of WikiText Results
We expand our perplexity evaluations to showcase additional benchmark datasets beyond WikiText, specifically C4 and the Pile, consistent with prior work. The results are summarized in Tab 13.
Model Dataset Baseline Wanda / +TRIM
Qwen2.5-7B
(65%)
WikiText 6.85 25.81 / 22.99
C4 11.88 34.64 / 31.15
Pile 7.79 22.61 / 20.39
LLaMA2-7B
(60%)
WikiText 5.47 10.77 / 10.64
C4 7.26 13.94 / 13.80
Pile 5.79 12.40 / 10.70
LLaMA2-13B
(65%)
WikiText 4.88 14.16 / 13.85
C4 6.73 19.32 / 18.98
Pile 5.44 15.42 / 15.15
Table 13: Comparison of baseline, Wanda, and Wanda+TRIM performance across different datasets.
We find that wikitext serves as a reliable indicator of overall perplexity behavior. In most cases, reductions in WikiText perplexity are accompanied by corresponding improvements on C4 and the Pile, affirming the broader effectiveness of our method beyond a single benchmark.
Appendix G Sensitivity Analysis of k
The hyperparameter k controls the number of iterations in the TRIM algorithm. A larger k allows for a more wide-ranging exploration of dimension-wise sparsity. Our findings indicate that k is a robust hyperparameter, with its impact depending on the difficulty of the pruning task. In challenging scenarios (e.g., Qwen2.5-14B at 70%), the iterative process is key, with performance improving significantly as k approaches our default of 10 before stabilizing. Conversely, in more straightforward scenarios (e.g., OPT-7B at 60%), the model’s performance is completely insensitive to k.
Table 14: Effect of different values for hyperparameter $k$ in TRIM. Perplexity is reported on Wikitext, averaged across two runs. Our default $k=10$ is underscored.
$k$ Qwen2.5-14B @ 70% OPT-7B @ 60% 3 99.46 14.73 6 66.70 14.73 10 56.67 14.73 15 58.60 14.73 25 59.98 14.73
Appendix H Hyperparameters
We publish the hyperparameters used for the layer-wise sparsity allocation methods. For OWL (Yin et al., 2024) we searched in the intervals $\lambda\in\{0.02,0.05,0.08,0.12,0.15,0.2\}$ and $\mathbf{M}\in\{3,5,7,10\}$. For AlphaPruning (Lu et al., 2024) we searched in the interval $\tau\in\{0.1,0.2,0.3,0.4\}$. All sweeps were done at 70% sparsity with one seed and Wanda as the underlying (Sun et al., 2024) pruning metric.
Table 15: Hyperparameters for the layerwise sparsity allocation methods that produce the results in this paper.
Model OWL (M, $\lambda$) AlphaPruning ($\tau$) OPT-6.7B (10, 12%) 0.1 OPT-13B (10, 5%) 0.1 LLaMA-2-7B (5, 20%) 0.3 LLaMA-2-13B (5, 15%) 0.3 Qwen2.5-3B (5, 12%) 0.1 Qwen2.5-7B (5, 12%) 0.1 Qwen2.5-14B (3, 20%) 0.3 Qwen2.5-32B (5, 15%) 0.3 Qwen2.5-72B (7, 8%) 0.1
Appendix I Pruning Metric Distribution Plots
We visualize the heterogeneity in the signal distribution at the dimension level in Fig 4.
Figure 4: Histogram showing the Gini coefficients for all layers of Qwen2.5-14B (K-proj). A high Gini means that the pruning metric (here: Wanda) is concentrated in fewer weights.
Appendix J Licensing
All used code, datasets and tasks are open-licensed under the MIT or CC by 4.0 license.
Appendix K Task Results
Full 0-shot task result tables (16, 17).
Table 16: Accuracies (%) for 7 zero-shot tasks for various models with unstructured 70% sparsity and based on Wanda. Higher is better. We add TRIM to non-uniform layer-wise allocation methods OWL and AlphaPruning.
Model Method BoolQ RTE Hella- Wino- ARC-e ARC-c OBQA Mean Swag Grande OPT-6.7B Dense 66.09 55.23 50.48 65.19 65.61 30.46 27.60 51.52 Alpha 62.14 55.96 28.99 52.49 40.07 17.06 12.20 38.41 +TRIM 62.17 53.43 35.72 54.54 48.86 22.35 18.80 42.27 OWL 62.17 51.99 32.90 53.35 47.10 19.96 17.20 40.67 +TRIM 62.20 52.71 33.13 52.80 48.27 20.82 18.20 41.16 OPT-13B Dense 65.87 57.76 52.46 65.19 67.09 32.85 27.00 52.60 Alpha 54.68 58.12 31.20 54.14 37.50 22.18 17.60 39.35 +TRIM 62.29 53.07 37.37 57.30 46.25 23.21 16.40 42.27 OWL 62.23 57.40 37.65 58.01 47.35 25.34 17.20 43.60 +TRIM 62.26 54.51 37.33 58.56 48.23 23.72 17.20 43.12 LLaMA-2-7B Dense 77.71 62.82 57.19 69.22 76.35 43.43 31.40 59.73 Alpha 62.26 52.71 34.74 61.64 43.22 22.10 17.40 42.01 +TRIM 62.45 53.07 34.74 58.56 42.97 22.18 17.20 41.60 OWL 62.35 52.71 36.91 60.93 48.11 22.87 19.20 43.30 +TRIM 62.39 52.71 37.15 60.22 47.81 23.46 19.80 43.36 LLaMA-2-13B Dense 80.55 65.34 60.05 72.14 79.42 48.46 35.20 63.02 Alpha 62.91 52.71 40.53 66.69 56.27 29.61 23.00 47.39 +TRIM 65.41 53.07 40.79 66.46 57.28 29.44 23.00 47.92 OWL 70.43 52.71 41.27 65.90 58.96 29.61 24.20 49.01 +TRIM 71.62 52.71 41.82 67.48 58.75 29.61 24.80 49.54 Qwen2.5-3B Dense 77.34 75.45 54.95 68.35 77.36 44.88 29.60 61.13 Alpha 53.91 53.43 27.72 50.04 33.33 18.17 11.40 35.43 +TRIM 61.71 52.71 27.68 50.20 34.47 17.66 12.80 36.75 OWL 62.17 52.71 27.97 49.96 33.42 17.49 12.00 36.53 +TRIM 62.17 52.71 28.00 48.70 36.07 16.89 12.00 36.65 Qwen2.5-7B Dense 84.68 81.59 60.01 73.01 80.47 47.78 33.20 65.82 Alpha 62.29 52.71 30.42 54.78 44.23 19.11 14.00 39.65 +TRIM 62.08 52.71 30.76 52.25 45.83 19.62 15.00 39.75 OWL 62.14 52.71 30.35 52.33 45.08 17.92 14.00 39.22 +TRIM 62.20 52.71 30.63 53.75 45.92 19.54 14.00 39.82 Qwen2.5-14B Dense 85.23 80.14 63.36 75.30 82.45 55.97 34.40 68.12 Alpha 62.17 52.71 29.86 54.46 47.09 18.94 14.20 39.92 +TRIM 62.20 52.71 30.66 55.56 48.57 20.22 15.20 40.73 OWL 62.17 52.07 31.82 57.14 48.36 19.28 15.60 40.92 +TRIM 62.17 52.71 32.54 58.88 50.42 19.45 15.60 41.68 Qwen2.5-32B Dense 87.12 81.59 64.98 75.22 80.85 53.07 33.80 68.09 Alpha 68.93 64.98 43.43 69.69 66.58 33.02 24.20 52.98 +TRIM 70.24 62.09 44.19 69.38 67.55 33.79 24.80 53.15 OWL 71.38 62.81 44.16 70.09 71.21 36.43 25.80 54.55 +TRIM 72.57 62.45 45.15 70.64 72.35 36.95 25.20 55.04 Qwen2.5-72B Dense 89.24 77.26 67.58 77.58 85.06 58.36 35.80 70.12 Alpha 77.43 79.06 49.83 74.35 74.12 43.00 27.80 60.80 +TRIM 80.31 80.87 50.09 74.59 74.66 43.17 28.60 61.76 OWL 76.61 78.34 50.77 74.43 76.26 44.88 27.80 61.30 +TRIM 81.80 80.50 50.92 75.06 76.26 44.11 28.60 62.47
Table 17: Accuracies (%) for 7 zero-shot tasks for various models with unstructured 80% sparsity and based on Wanda. Higher is better. We add TRIM to non-uniform layer-wise allocation methods OWL and AlphaPruning.
Model Method BoolQ RTE Hella- Wino- ARC-e ARC-c OBQA Mean Swag Grande OPT-6.7B Dense 66.09 55.23 50.48 65.19 65.61 30.46 27.60 51.52 Alpha 41.04 52.71 25.93 50.20 27.44 19.97 11.60 32.70 +TRIM 60.70 50.18 26.57 51.38 31.19 18.69 11.00 35.67 OWL 38.62 51.99 25.90 50.75 25.97 20.65 14.40 32.61 +TRIM 61.93 52.71 26.86 51.38 33.00 17.41 10.80 36.30 OPT-13B Dense 65.87 57.76 52.46 65.19 67.09 32.85 27.00 52.60 Alpha 60.83 51.26 25.83 49.49 26.94 19.96 11.60 35.13 +TRIM 61.87 52.71 26.75 52.25 29.80 19.54 13.00 36.56 OWL 56.39 52.71 26.10 51.14 26.89 20.90 12.40 35.22 +TRIM 58.01 52.71 26.99 50.67 30.35 17.75 11.80 35.47 LLaMA-2-7B Dense 77.71 62.82 57.19 69.22 76.35 43.43 31.40 59.73 Alpha 37.89 52.71 26.20 47.28 26.73 20.14 12.60 31.94 +TRIM 37.89 52.71 26.18 47.04 26.56 20.56 13.00 31.99 OWL 38.19 52.71 26.64 50.36 27.82 18.69 13.80 32.60 +TRIM 38.26 52.71 26.50 51.38 27.86 19.45 13.40 32.79 LLaMA-2-13B Dense 80.55 65.34 60.05 72.14 79.42 48.46 35.20 63.02 Alpha 62.05 52.71 27.74 52.72 30.68 19.28 11.80 36.71 +TRIM 62.17 52.71 27.90 51.62 30.85 19.88 12.00 36.73 OWL 61.41 52.71 27.49 50.75 30.35 19.11 11.80 36.23 +TRIM 62.08 52.71 28.06 53.04 30.64 19.37 12.20 36.87 Qwen2.5-3B Dense 77.34 75.45 54.95 68.35 77.36 44.88 29.60 61.13 Alpha 37.77 55.23 26.12 50.04 27.06 19.80 13.40 32.77 +TRIM 44.28 49.46 26.47 48.54 28.58 20.56 14.20 33.16 OWL 56.73 52.71 26.78 51.22 30.30 19.03 11.20 35.42 +TRIM 54.10 52.71 26.81 49.25 29.71 18.17 13.00 34.82 Qwen2.5-7B Dense 84.68 81.59 60.01 73.01 80.47 47.78 33.20 65.82 Alpha 37.92 52.71 26.91 49.88 28.37 19.03 14.40 32.74 +TRIM 37.77 53.06 27.21 49.96 30.01 18.86 14.00 32.98 OWL 37.89 52.71 27.14 47.28 30.51 17.83 12.80 32.31 +TRIM 38.29 52.71 27.30 47.83 30.09 18.09 14.20 32.64 Qwen2.5-14B Dense 85.23 80.14 63.36 75.30 82.45 55.97 34.40 68.12 Alpha 59.17 52.71 26.94 48.93 30.43 16.64 12.40 35.32 +TRIM 50.67 52.71 26.86 50.12 30.89 16.55 13.00 34.40 OWL 44.59 52.71 27.27 49.17 31.94 16.64 13.60 33.70 +TRIM 42.08 52.71 27.08 49.25 32.15 16.98 13.00 33.32 Qwen2.5-32B Dense 87.12 81.59 64.98 75.22 80.85 53.07 33.80 68.09 Alpha 59.14 52.71 27.61 48.22 35.19 17.15 12.40 36.06 +TRIM 61.83 53.43 28.79 49.49 37.46 17.49 13.60 37.44 OWL 62.08 52.71 28.49 49.57 38.38 17.24 13.00 37.35 +TRIM 62.20 52.71 28.35 50.20 39.18 16.30 13.40 37.48 Qwen2.5-72B Dense 89.24 77.26 67.58 77.58 85.06 58.36 35.80 70.12 Alpha 62.17 52.71 32.14 58.41 41.12 21.50 14.20 40.32 +TRIM 62.17 52.71 32.50 58.72 41.24 21.16 15.60 40.59 OWL 62.17 52.71 32.63 58.88 43.56 21.67 16.20 41.12 +TRIM 62.17 52.71 32.69 57.85 44.28 21.42 16.00 41.02